### v1

Abstract
While the auto-bidding literature predominantly considers independent bidding, we investigate the coordination problem among multiple auto-bidders in online advertising platforms. Two motivating scenarios are: collaborative bidding among multiple distinct bidders managed by a third-party bidding agent, and strategic bid selection for multiple ad campaigns managed by a single advertiser. We formalize this coordination problem as a theoretical model and demonstrate that a straightforward coordination mechanism, where only the highest-value bidder competes with outside bids, strictly dominates independent bidding, improving both Return-on-Spend (RoS) compliance and the total value accrued for each participating auto-bidder or ad campaign. Additionally, our simulations on synthetic and real-world datasets support the theoretical result that coordinated mechanism outperforms independent bidding. These findings highlight both the theoretical potential and the practical robustness of coordination in auto-bidding in online auctions.
1 Introduction
Online advertising has been demonstrated as a prominent approach to monetize search engines, news feeds, social media, and e-commerce. Among them, real-time auctions have been used widely to connect advertisers and users efficiently. In the modern online advertising platforms, advertisers endeavor to maximize their campaign effectiveness, quantified by conversions or other pertinent metrics, through precise bid management that considers targeted return-on-spend (RoS) [aggarwal2019autobidding], which is also well-known as auto-bidding problems. To this end, a wide spectrum of bidding methodologies has been formulated, leveraging concepts from optimization theory, online learning paradigms, and game-theoretic principles [zhao2018deep, aggarwal2019autobidding, lee2013real, babaioff2021non, golrezaei2021bidding, deng2021towards, balseiro2021landscape, gao2022bidding, stram2024mystique].
Despite the wealth of sophisticated research in online advertising, existing literature predominantly concentrates on optimizing the value or utility from the perspective of individual bidders. This focus, while valuable, often overlooks a crucial aspect of real-world advertising dynamics: the potential for coordinated bidding strategies. In practice, it is common for multiple bidders to form coalitions, perhaps by authorizing a third-party advertising platform to manage their bids collectively. Alternatively, a single large advertiser, such as major e-commerce players like Amazon, Temu, or Shein, might control numerous distinct advertising campaigns. These entities often have a portfolio of advertisements they are willing to display to users, necessitating a coordinated approach across their various campaigns to maximize overall effectiveness, rather than optimizing each in isolation. Consequently, a deeper understanding of how auto-bidding algorithms can be coordinated across multiple allied bidders or campaigns, and what are the benefits of coordination over independent bidding, is becoming increasingly critical. Such understandings might be useful for industry practices and unlock further efficiencies in online advertising markets.
In this work, we investigate the coordination problem among multiple auto-bidding algorithms in repeated auctions. We formulate a theoretical model in which $N$ auto-bidders participate in $T$ rounds of second-price auctions. These bidders may form a coalition to determine their bids jointly, competing against bidders outside the coalition. Each auto-bidder is modeled as a value-maximizing agent subject to a Return on Spend (RoS) constraint (i.e., non-negative overall utility). We consider a simple form of coordination: only the bidder with the highest value inside the coalition submits a positive bid to compete against the outside bidders, while the other N1N-1 coalition members refrain from competing. In contrast, in the independent bidding scenario, all the $N$ bidders submit positive bids to compete against each other and the outside bidders. We aim to compare the coalition bidders’ welfare in the coordinated and independent bidding scenarios.
The main contributions of this paper are as follows. Theoretically, we prove that the highest-value-bidder-compete coordination mechanism mentioned above, although straightforward, enjoys a significant advantage over independent bidding: By coordination, the bidders in the coalition acquire higher values, as well as achieve lower RoS constraint violations, compared to independent bidding.
Specifically, we prove the following two results: (1) As long as the coalition bidders overbid (a common feature for value-maximizing bidding algorithms), then no matter what auto-bidding algorithms they use, coordination always increases the total utility (value - payment) of each bidder in the coalition over $T$ rounds, compared to the independent bidding scenario. This means that the bidders’ RoS constraints are better satisfied. (2) If coalition bidders run a state-of-the-art value-maximizing algorithm (a mirror-descent algorithm in feng2023online), then the total value obtained by the coalition is higher under coordination than under independent bidding.
Our theoretical results are proven under two mild assumptions. The first one (Assumption 1) is that the highest and second-highest values among the bidders inside the coalition should be competitive against the highest bid outside of the coalition. Without such an assumption, the coalition as a whole is too weak to compete against the outside bidders, and the problem becomes non-interesting. The second assumption is that the bidders in the coalition have independent and identically distributed (i.i.d.) valuations. This i.i.d. assumption is not only for theoretical simplicity but also reasonable in practice. For the two motivating scenarios mentioned above (multiple bidders forming a coalition and a single bidder controlling multiple advertising campaigns), the bidders/campaigns selected by the auctioneer to compete usually share similar features, justifying i.i.d. valuations. In addition, we show in Section 5 that our analysis can be extended to non-i.i.d. settings.
Finally, (in Section 6) we demonstrate by experiments, on both synthetic and real-world datasets, that our coordinated mechanism consistently outperforms independent bidding, validating our theoretical results under realistic market conditions.
1.1 Related Work
Bidder coordination, often referred to as cartels or collusion in economics, has been extensively studied in traditional auction theory [robinson1985collusion, hendricks1989collusion, marshall2007bidder]. Prior research has studied collusive behavior in first-price auctions [lopomo2011bidder, pesendorfer2000study], second-price auctions [mailath1991collusion, graham1987collusive], collusion-proof mechanism design [che2009optimal], auction design under collusion [pavlov2008auction], and collusion detection [chotibhongs2012analysis]. While these studies typically consider bidders forming collusions autonomously and reaching equilibrium outcomes in static auction settings [bergemann2016bayes, fu2025learning], we instead introduce a central planner that coordinates bidders to improve their total welfare in dynamic auction environments.
Recently, coordination behavior has also been observed in online advertising auctions [decarolis_bid_2023], sparking growing interest in understanding and analyzing such phenomena. decarolis2020marketing studied the impact of coordination on the revenue and efficiency of GSP and VCG auctions. romano2022power investigated the computational challenges faced by a media agency coordinating bidders under those mechanisms. chen_coordinated_2023 analyzed coordinated online bidding in repeated second-price auctions with budget constraints. While those works focus on coordinating utility maximizers, we study value maximizers. To the best of our knowledge, this is the first work to investigate coordination among value maximizers, which exhibit different properties from utility maximizers.
Our work is also related to the large literature on auto-bidding. For example, paes2024complex studied the dynamics of systems with multiple auto-bidders and revealed complex behaviors such as bi-stability, periodic orbits, and quasi-periodicity, while aggarwal2025multi analyzed optimal bidding strategies for advertisers across multiple platforms. For a comprehensive overview of the literature, we refer readers to the survey by aggarwal_auto-bidding_2024. Our study complements the predominant individual-optimization perspective of the auto-bidding literature.
2 Model: Coordinated Auto-Bidding
Repeated Second-Price Auctions with Coalition.
We consider the scenario where $N$ auto-bidders (or bidders) participate in $T$ rounds of repeated second-price auctions. In each round $t\in[T]=\{1,\ldots,T\}$, one item (e.g., ad slot) is for sale, and each bidder $i\in[N]=\{1,\ldots,N\}$ has a value $v_{i,t}\in[0,B]$ for the item i.i.d. sampled from a continuous distribution $F$. Let $\bm{v}_{t}=(v_{1,t},\ldots,v_{N,t})$ be the vector of values of all bidders at round $t$. Let $b_{i,t}$ denote bidder $i$’s bid at round $t$. The $N$ bidders may form a coalition. Our model captures two possible types of coalitions in ad auctions:
  1. 1.
    NN advertisers, each with an ad, instruct their auto-bidding algorithms to coordinate with each other before bidding.
  2. 2.
    One advertiser owns $N$ ads. Instead of using $N$ auto-bidding algorithms to bid for the $N$ ads independently, the advertiser instructs the algorithms to coordinate.
In either case, usually in practice, only similar advertisers/ads are selected to compete for one ad slot, which justifies the i.i.d. value assumption. In addition, we will show in Section 5 that our results extend to non-i.i.d. value settings.
There is a competing bid $d_{t}^{O}\in[0,B]$ from outside of the coalition, assumed to be i.i.d. sampled from another distribution $D$ every round. Our model also captures second-price auctions with reservation price, because $d_{t}^{O}$ can be the maximum between the reservation price and the outside competing bid. We denote the competing bid faced by bidder $i$ at round $t$ as
$d_{i,t}=\max\big\{d^{O}_{t},\max_{j\in[N]\setminus\{i\}}b_{j,t}\,\big\}.$ (1)
Let $x_{i,t}:=\mathbb{I}\{b_{i,t}\geq d_{i,t}\}\in\{0,1\}$ indicate whether bidder $i$ wins the item (ignoring tie-breaking for now), so bidder $i$’s payment is $p_{i,t}:=x_{i,t}d_{i,t}$ and utility is $u_{i,t}:=x_{i,t}(v_{i,t}-d_{i,t})$ at round $t$.
Value Maximization under RoS Constraint.
Each auto-bidder aims to maximize its total value while adhering to the Return on Spend (RoS) constraint. The formal definition of the optimization problem for bidder $i$ is as follows:
$\max_{\{b_{i,t}\}_{t=1}^{T}}$ $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}$ (bidder $i$’s value)
s.t. $\sum_{t=1}^{T}u_{i,t}=\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\sum_{t=1}^{T}p_{i,t}\geq 0$ (RoS constraint)
The RoS constraint ensures that for every dollar spent, at least one dollar of value is generated; it is equivalent to requiring the bidder’s total utility to be non-negative. [^1] In fact, the general RoS constraint requires $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\tau\cdot\sum_{t=1}^{T}p_{i,t}\geq 0$, i.e., at least $\tau$ dollar of value will be generated for every dollar spend, however, it is without loss of generality to assume $\tau=1$ in the above optimization problem because we can rescale the bidders’ values by dividing $\tau$.
Independent and Coordinated Auto-Bidding
Each auto-bidder runs an algorithm $A$ that learns to bid from history, in order to maximize the bidder’s total value subject to RoS constraint. Formally, at each round $t$, bidder $i$’s algorithm $A(\cdot)$ takes the historical information $H_{i,t}$ available to bidder $i$ (for example, the past values, bids, and allocations of the bidder: $H_{i,t}=\{v_{i,t^{\prime}},b_{i,t^{\prime}},x_{i,t^{\prime}}\}_{t^{\prime}=1}^{t-1}\cup\{v_{i,t}\}$) as input and outputs the bid $b_{i,t}$ for the current round. We refer to the scenario where the auto-bidding algorithms are run independently as independent bidding, formalized in Algorithm 1.
for $t=1,2,\dots,T$ do
   Each bidder $i\in[N]$ observes value $v_{i,t}$ and bids $b_{i,t}=A(H_{i,t})$;
   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
end for
ALGORITHM 1 Independent Bidding
Next, we define a coordinated bidding scenario where, at each round, only the highest-value bidder in the coalition places a positive bid while all other bidders bid zero. We assume that ties are resolved naturally. [^2] Ties happen with probability 0 given a continuous value distribution $F$. This is formalized in Algorithm 2.
for $t=1,2,\dots,T$ do
   Each bidder $i\in[N]$ observes value $v_{i,t}$;
   Let $i^{*}=\arg\max_{i\in[N]}v_{i,t}$;
   Bidder $i^{*}$ bids $b_{i^{*},t}=A(H_{i^{*},t})$; all $i\neq i^{*}$ bid $b_{i,t}=0$;
   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
end for
ALGORITHM 2 Coordinated Bidding
3 Coordination Reduces RoS Violation
In this section, we analyze the extent to which the RoS constraint of every auto-bidder is maintained/violated. We will show that, under mild assumptions, coordinated bidding results in a smaller RoS violation compared to independent bidding. Equivalently, every bidder in the coalition obtains a higher utility through coordination.
Formally, let $U_{i}^{C,A},U_{i}^{I,A}$ be the total utility of bidder $i$ during the $T$ rounds using auto-bidding algorithm $A$ under coordinated and independent bidding scenarios, respectively:
$U_{i}^{C,A}=\sum_{t=1}^{T}u_{i,t}^{C,A},\quad U_{i}^{I,A}=\sum_{t=1}^{T}u_{i,t}^{I,A}.$ (2)
RoS constraint requires $U_{i}^{C,A}\geq 0$ and $U_{i}^{I,A}\geq 0$. If $U_{i}^{C,A}\geq-c$ or $U_{i}^{I,A}\geq-c$, we say the RoS constraint for bidder $i$ is violated by $c$.
We allow the bidders to use any auto-bidding algorithm that (weakly) overbids: namely, $A(\cdot)$ can be an arbitrary function that satisfies $b_{i,t}=A(H_{i,t})\geq v_{i,t}$. Note that we are considering value-maximizing bidders in second-price auctions. If the bidders underbid ($b_{i,t}<v_{i,t}$), they will always obtain non-negative utility and satisfy the RoS constraint, but not necessarily maximize their total values. So we assume the bidders to overbid.
We make a reasonable assumption on the distributions of value and outside bid, $F$ and $D$. Let $(x)_{+}=\max\{x,0\}$ denote the positive part of a number. Let $v_{(N)},v_{(N-1)}$ be the largest and second-largest values among $N$ i.i.d. samples from $F$. We assume that: in expectation, the advantage of the second-largest value $v_{(N-1)}$ over the outside bid $d^{O}$ is larger than the advantage of $d^{O}$ over the largest value $v_{(N)}$:
Assumption 1
$\mathbb{E}_{F,D}\big[\big(v_{(N-1)}-d^{O}\big)_{+}-\big(d^{O}-v_{(N)}\big)_{+}\big]\geq\Delta>0$.
Intuitively, Assumption 1 says that the largest two values among the coalition bidders should be competitive enough against the outside bid. With large $v_{(N-1)}$ and $v_{(N)}$, $\mathbb{E}[(v_{(N-1)}-d^{O})_{+}]$ is large while $\mathbb{E}[(d^{O}-v_{(N)})_{+}]$ is small, which satisfy the assumption.
Example 1
The following examples satisfy Assumption 1:
  • $N=4,F=U[0,1],D=U[0,1]$, with $\Delta=1/6$.
  • $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$, with $\Delta=1/40$.
We will perform simulations on these examples in Section 6.
The main result of this section is the following theorem, which shows that coordinated bidding strictly improves each bidder’s total utility compared to independent bidding. This directly implies that each bidder’s RoS constraint violation under coordination is smaller.
Theorem 3.1(Coordination reduces RoS violation)
Assume Assumption 1. For any auto-bidding algorithm $A$ that overbids, for every bidder $i\in[N]$, with probability at least $1-\exp(-\frac{\Delta^{2}T}{32B^{2}N^{2}})$,
$U_{i}^{C,A}~\geq~U_{i}^{I,A}+\tfrac{\Delta T}{2N}.$
The proof of Theorem 3.1 involves an interesting comparison between the coordinated bidding scenario and the truthful bidding scenario where all bidders bid their values exactly, and a comparison between truthful bidding and independent auto-bidding. These two comparisons together imply that coordinated bidding generates higher utility for every coalition member than independent bidding. We present the detailed proof below.
3.1 Proof of Theorem 3.1
Let $U_{i}^{\mathrm{Truth}}=\sum_{t=1}^{T}u_{i,t}^{\mathrm{Truth}}$ be the total utility of bidder $i$ when all bidders in the coalition bid their values truthfully: $b^{\mathrm{Truth}}_{i,t}=v_{i,t}$. We will prove Theorem 3.1 by proving $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$ and $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$. The second inequality is formalized below:
Lemma 1
For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$.
Proof
This lemma directly follows from the dominant-strategy-incentive-compatibility of second-price auction, which ensures that each bidder’s truthful-bidding utility $u^{\mathrm{Truth}}_{i,t}$ is weakly larger than the non-truthful bidding utility $u^{I,A}_{i,t}$ at every round. ∎
We then prove $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$:
Lemma 2
[…]
Proof
For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, consider the difference between bidder $i$’s coordinated auto-bidding utility and truthful bidding utility at round $t$:
$\Delta_{i,t}$ $=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$
$=(v_{i,t}-d_{t}^{O})\cdot\mathbb{I}[b_{i,t}^{C,A}\geq d_{t}^{O}]\cdot\mathbb{I}[v_{i,t}=\max_{j\in[N]}v_{j,t}]$
$\quad-(v_{i,t}-d_{i,t}^{\mathrm{Truth}})\cdot\mathbb{I}[v_{i,t}\geq d_{i,t}^{\mathrm{Truth}}]$
where $d_{i,T}^{\mathrm{Truth}}=\max\{d_{t}^{O},\max_{j\neq i}v_{j,t}\}$ is the competing bid for bidder $i$ under truthful bidding. We do a case analysis:
  • If $v_{i,t}\neq\max_{j\in[N]}v_{j,t}$, then bidder $i$ loses under both coordinated bidding and truthful bidding, so $u_{i,t}^{C,A}=u_{i,t}^{\mathrm{Truth}}=0$ and $\Delta_{i,t}=0$.
  • If $v_{i,t}=\max_{j\in[N]}v_{j,t}$, then
    $u_{i,t}^{C,A}=(v_{i,t}-d_{t}^{O})\cdot\mathbb{I}[b_{i,t}^{C,A}\geq d_{t}^{O}]$
    $u_{i,t}^{\mathrm{Truth}}=(v_{i,t}-d_{t}^{\mathrm{Truth}})\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}].$
    Because the auto-bidder is overbidding ($v_{i,t}\leq b_{i,t}^{C,A}$), there are three further cases:
    • $d_{t}^{O}<v_{i,t}\leq b_{i,t}^{C,A}$: in this case, $\Delta_{i,t}=d_{t}^{\mathrm{Truth}}-d_{t}^{O}\geq 0$.
    • $v_{i,t}\leq b_{i,t}^{C,A}<d_{t}^{O}$: in this case, the bidder loses under both scenarios, so $\Delta_{i,t}=0$.
    • $v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}$: in this case, the bidder loses under truthful bidding but wins under coordinated bidding, so $\Delta_{i,t}=u_{i,t}^{C,A}=v_{i,t}-d_{t}^{O}\leq 0$.
Then, we condition on any history $H_{t-1}$, randomize over $\bm{v}_{t}$ and $d_{t}^{O}$, and analyze the expectation of $\Delta_{i,t}$. Let $F_{(N)}$ be the distribution of the largest value $v_{(N)}$ among $N$ samples from $F$.
[…]
$=\frac{1}{N}\bigg(\Pr_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\Big[v_{i,t}>d_{t}^{O}\Big]\cdot\mathbb{E}\Big[d_{t}^{\mathrm{Truth}}-d_{t}^{O}\,\big|\,v_{i,t}\geq d_{t}^{O}\Big]$
$\quad+\Pr_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\Big[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\cdot\mathbb{E}\Big[v_{i,t}-d_{t}^{O}\,\big|\,v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\bigg)$
$=\frac{1}{N}\bigg(\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}>d_{t}^{O}\big]\cdot\big(\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\big)\Big]$
$\qquad+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg).$
Because $\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}>0$ implies $\mathbb{I}\big[v_{(N)}>d_{t}^{O}\big]=1$, the above is equal to
$=\frac{1}{N}\bigg(\mathbb{E}\Big[\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\Big]$
$\qquad+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
$\geq\frac{1}{N}\bigg(\mathbb{E}\Big[\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\Big]+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
$=\frac{1}{N}\bigg(\mathbb{E}\Big[\big(v_{(N-1)}-d_{t}^{O}\big)_{+}\Big]-\mathbb{E}\Big[\big(d_{t}^{O}-v_{(N)}\big)_{+}\Big]\bigg).$
Using Assumption 1, we obtain
[…]
Because values and outside bids are bounded by $B$, $\Delta_{i,t}$ is bounded by [B,B][-B,B]. So, by Azuma’s inequality,
[…]
Letting ${\varepsilon}=\frac{\Delta T}{2N}$ proves the lemma. ∎
Lemmas 1 and 2 together prove Theorem 3.1.
4 Coordination Increases Bidders’ Value
In this section, we show that coordination can increase the total value of auto-bidders, compared to independent bidding.
We will focus on a specific auto-bidding algorithm: the mirror-descent algorithm in feng2023online. It is the state-of-the-art algorithm for value maximization under RoS constraint, with near-optimal performances for the bidders, achieving $O(\sqrt{T})$ regrets. The algorithm maintains a parameter $\lambda_{i,t}$ for each bidder (which corresponds to the Lagrange multiplier of the dual optimization problem), so the bidder overbids by $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$ at each round. After observing a utility feedback, the bidder adjusts the parameter $\lambda_{i,t}$ in the opposite direction of the utility (i.e., positive utility decreases $\lambda_{i,t}$). The full algorithm is given in Algorithm 3.
Initialize multiplier $\lambda_{i,1}=1$, learning rate $\alpha=1/\sqrt{T}$;
for $t=1,2,\dots,T$ do
   Observe value $v_{i,t}$ and set bid $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$;
   Observe allocation $x_{i,t}$ and payment $p_{i,t}$, compute utility $g_{i,t}=v_{i,t}\cdot x_{i,t}-p_{i,t}$;
   Update the multiplier: $\lambda_{i,t+1}=\lambda_{i,t}\exp(-\alpha g_{i,t})$;
end for
ALGORITHM 3 Auto-bidding algorithm in feng2023online
For Algorithm 3, we show that coordinated auto-bidders obtain higher total value than independent auto-bidders. Let
$V_{t}^{C}=\sum_{i=1}^{N}v_{i,t}x^{C}_{i,t},\quad V_{t}^{I}=\sum_{i=1}^{N}v_{i,t}x^{I}_{i,t}$
be the total values obtained by the $N$ bidders at round $t$ in the coordinated case and independent case, respectively.
Theorem 4.1(Coordination Increases Total Value)
Assume 1, and suppose bidders run Algorithm 3. As $T\to\infty$, in expectation, the bidders’ total value under coordinated bidding is larger than that under independent bidding:
$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0.$
To prove Theorem 4.1, we first reduce the $N$-bidder coordinated scenario to a single-bidder problem where the single bidder’s value is the maximum among the coalition members. Using this reduction, we prove an interesting fact: under Assumption 1, the multipliers $\lambda_{i,t}^{C}$ of coordinated bidders converge to 0 in the limit as $t\to\infty$, meaning that the coordinated bidders will eventually submit arbitrarily high bids. This implies that the coordinated coalition will obtain a higher value than the independent case. We present the full proof below.
4.1 Proof of Theorem 4.1
Reduction to a single-bidder scenario.
To prove Theorem 4.1, we first draw a connection between the $N$-bidder coordination scenario and the scenario where a single bidder, whose value equals the highest value among the $N$ bidders, competes against the outside bid.
Formally, let $v_{(N)}=\max\{v_{1},\ldots,v_{N}\}$ denote the highest value among $N$ i.i.d. samples from $F$. Let $F_{(N)}$ be the distribution of $v_{(N)}$. Let $G(\lambda)$ be the expected utility of a single bidder with value $v_{(N)}\sim F_{(N)}$ and bidding using multiplier $\lambda>0$, competing against the outside bid $d^{O}\sim D$:
$G(\lambda)=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Big[(v_{(N)}-d^{O})\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
Let $G_{i}^{C}(\lambda_{i})$ be the expected utility of bidder $i$ with value $v_{i}\sim F$ and multiplier $\lambda_{i}>0$ in the $N$-bidder coordinated scenario:
$G_{i}^{C}(\lambda_{i})=\mathbb{E}_{v_{1},\ldots,v_{N}\sim F,d^{O}\sim D}\big[(v_{i}-d^{O})x_{i}(\bm{v},\lambda_{i})\big]$
where $x_{i}(\bm{v},\lambda_{i})=1$ if $v_{i}=\max_{j\in[N]}\{v_{j}\}$ and $(1+1/\lambda_{i})v_{i}>d^{O}$, and 0 otherwise.
We show that the $N$-bidder coordinated scenario is “equivalent” to the single-bidder scenario in the following sense:
Observation 1
$G_{i}^{C}(\lambda_{i})=\frac{1}{N}G(\lambda_{i})$.
Proof
In the coordinated scenario, bidder $i$ obtains utility 0 if its value is not the highest among the $N$ bidders. The probability that bidder $i$ has the highest value is $1/N$ because the $N$ bidders’ values are i.i.d. So,
$G_{i}^{C}(\lambda_{i})$ $=0+\tfrac{1}{N}\mathbb{E}_{v_{i}\sim F|v_{i}=v_{(N)}}\big[(v_{i}-d^{O})\mathbb{I}[(1+1/\lambda_{i})v_{i}>d^{O}]\big]$
$=\tfrac{1}{N}\mathbb{E}_{v_{(N)}\sim F_{(N)}}\big[(v_{(N)}-d^{O})\mathbb{I}[(1+1/\lambda_{i})v_{(N)}>d^{O}]\big]$
$=\tfrac{1}{N}G(\lambda_{i}).$
Since the $N$-bidder coordinated scenario is equivalent to a single-bidder scenario, we can now analyze the single-bidder scenario. We present two useful lemmas regarding the expected utility function $G(\lambda)$ of the single bidder.
Lemma 3
$G(\lambda)$ is monotonically increasing in $\lambda$.
Proof
See Appendix 0.A. ∎
Lemma 4
Under Assumption 1, for any $\lambda>0$, $G(\lambda)\geq\Delta>0$.
Proof
By Lemma 3, $G(\lambda)$ is increasing in $\lambda$. So,
$G(\lambda)$ $\geq\lim_{\mu\to 0^{+}}G(\mu)$
$=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\big[v_{(N)}-d^{O}\big]$
$=\mathbb{E}\big[(v_{(N)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]$
$\geq\mathbb{E}\big[(v_{(N-1)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]\geq\Delta>0$
where the last $\geq$” follows from Assumption 1. ∎
Intuitively, Lemma 4 shows that the expected utility obtained by the highest-value coalition bidder is always positive (by a margin of $\Delta>0$). This means that the bidder’s multiplier $\lambda_{t}$, following mirror descent updates, will decrease in expectation every round and eventually converge to 0. We formalize this intuition below.
Coordinated multipliers converge to 0.
Using the above characterizations, we show an interesting fact: the multipliers $\lambda_{i,t}^{C}$ of auto-bidders running Algorithm 3 in the coordinated scenario converge to 0 as $t\to\infty$. This means that the coalition will eventually submit arbitrarily high bids $b_{i,t}^{C}=(1+\frac{1}{\lambda_{i,t}^{C}})v_{i,t}\to\infty$.
Lemma 5
Under Assumption 1, for any $i\in[N],t\in[T]$, with probability at least $1-\exp\big(-\frac{\Delta^{2}t}{32B^{2}N^{2}}\big)$, $\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\big(-\frac{\Delta t}{2N\sqrt{T}}\big)$.
Proof
According to Algorithm 3, $\lambda_{i,t+1}^{C}$ is equal to
$\lambda_{i,t+1}^{C}=\lambda_{i,1}\exp\Big(-\alpha\sum_{t^{\prime}=1}^{t}g_{i,t^{\prime}}\Big)$
where each $g_{i,t^{\prime}}$ is an unbiased estimator of bidder $i$’s expected utility $G_{i}^{C}(\lambda_{i,t^{\prime}})=\mathbb{E}[g_{i,t^{\prime}}]$ at round $t^{\prime}$. If bidder $i$ is not the highest-value bidder in round $t^{\prime}$, then $g_{i,t^{\prime}}=0$, meaning that the bidder’s multiplier $\lambda_{i,t^{\prime}}^{C}$ is not updated in that round. According to Observation 1 and Lemma 4, bidder $i$’s expected utility is at least
$\mathbb{E}[g_{i,t^{\prime}}]=G_{i}^{C}(\lambda_{i,t^{\prime}})=\frac{1}{N}G(\lambda_{i,t^{\prime}})\geq\frac{\Delta}{N}.$
Because values and outside bids are bounded by $B$, the realized utility $g_{i,t^{\prime}}\in[-B,B]$. So, by Azuma’s inequality:
$\Pr\left[\sum_{t^{\prime}=1}^{t}g_{i,t^{\prime}}<t\cdot\frac{\Delta}{N}-{\varepsilon}\right]\leq\exp\Big(-\frac{{\varepsilon}^{2}}{8B^{2}t}\Big).$
Let ${\varepsilon}=\frac{t\Delta}{2N}$. We have with probability at least $1-\exp(-\frac{t\Delta^{2}}{32B^{2}N^{2}})$,
$\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\Big(-\alpha\big(t\cdot\frac{\Delta}{N}-{\varepsilon}\big)\Big)=\lambda_{i,1}\exp\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big).$
where we plugged in $\alpha=1/\sqrt{T}$. ∎
Coordinated total value is larger than independent total value.
Using the fact that coordinated multipliers converge to 0, we show that the total value obtained by coordinated auto-bidders is larger than that under independent bidding in the limit $T\to\infty$.
Let $V_{(N)}(\lambda)$ be the expected value obtained by a single bidder whose value is the highest value $v_{(N)}$ among $N$ samples from $F$ and who uses multiplier $\lambda>0$ to bid against outside bid $d^{O}\sim D$:
$V_{(N)}(\lambda)=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Big[v_{(N)}\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
Because $V_{(N)}(\lambda)$ is monotone in $\lambda$ and bounded, the limit
$V_{(N)}(0^{+})=\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$
exists. Note that the expected total value of $N$ coordinated bidders with multipliers $\lambda_{1}^{C},\ldots,\lambda_{N}^{C}$ is
$V^{C}(\lambda_{1}^{C},\ldots,\lambda_{N}^{C})=\mathbb{E}\Big[v_{i^{*}}\!\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda_{i^{*}}})v_{i^{*}}>d^{O}\big]\Big]=\mathbb{E}\big[V_{(N)}(\lambda_{i^{*}}^{C})\big]$
where $i^{*}=\operatorname*{arg\,max}_{i\in[N]}v_{i}$. This means that the expected average total value of the coordinated bidders is
$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{C}\Big]=\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i^{*}_{t},t}^{C})\Big]$ (3)
where $i^{*}_{t}\in\{1,\dots,N\}$ is the highest-value bidder in round $t$.
Let $V^{I}(\lambda_{1},\ldots,\lambda_{N})$ be the expected total value of $N$ independent bidders with multipliers $\lambda_{1},\ldots,\lambda_{N}>0$:
$V^{I}(\lambda_{1},\ldots,\lambda_{N})=\mathbb{E}_{\bm{v}\sim F,d^{O}\sim D}\Big[\sum_{i=1}^{N}v_{i}x_{i}(\bm{b}^{I})\Big]$
where $\bm{b}^{I}=\big((1+\tfrac{1}{\lambda_{1}})v_{1},\ldots,(1+\tfrac{1}{\lambda_{N}})v_{N},d^{O}\big)$.
Lemma 6
$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$, for any $\lambda_{1},\ldots,\lambda_{N}>0$.
Proof
Pick $\lambda_{\varepsilon}<\min\{\lambda_{1},\dots,\lambda_{N}\}$. If some bidder $i\in[N]$ wins the item under independent bidding, then their bid $b_{i}^{I}=(1+\frac{1}{\lambda_{i}})v_{i}$ is larger than the outside bid $d^{O}$. In that case, suppose the highest-value bidder uses multiplier $\lambda_{\varepsilon}$, then their bid is $(1+\frac{1}{\lambda_{\varepsilon}})v_{(N)}\geq(1+\frac{1}{\lambda_{i}})v_{i}=b_{i}^{I}>d^{O}$, so the highest-value bidder also wins in the single-bidder scenario. Note that the value obtained by the highest-value bidder is at least $v_{i}$, so the expected value $V_{(N)}(\lambda_{\varepsilon})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$. Letting $\lambda_{\varepsilon}\to 0$ proves the lemma. ∎
Proof(Proof of Theorem 4.1)
By Lemma 5, with constant $c\coloneqq\frac{\Delta^{2}}{32B^{2}N^{2}}$, for any fixed $i$ and $t\geq 1$, the coordinated multiplier $\lambda_{i,t+1}^{C}$ satisfies
$\Pr\!\left[\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\right]\;\geq\;1-e^{-ct}.$ (4)
Fix $\varepsilon>0$. Since $V_{(N)}(0^{+})$ exists, there is $\delta=\delta(\varepsilon)>0$ such that
$\big|V_{(N)}(0^{+})-V_{(N)}(\lambda)\big|<\varepsilon\quad\text{for all }\lambda\in(0,\delta].$ (5)
Define
$m_{T}\;\coloneqq\;\left\lceil\frac{2N\sqrt{T}}{\Delta}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\right\rceil~=~O\Big(\sqrt{T}\log\frac{1}{\delta}\Big),$
so for all $t\geq m_{T}$,
$\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\leq\lambda_{i,1}\exp\!\Big(-\frac{m_{T}\Delta}{2N\sqrt{T}}\Big)\leq\delta.$
Let $E_{i,t}$ be the event inside the probability in (4). A union bound over $i=1,\dots,N$ and $t\geq m_{T}$ gives
$\Pr\!\Big[\bigcap_{i=1}^{N}\bigcap_{t=m_{T}}^{T}E_{i,t}\Big]\;\geq\;1-NTe^{-cm_{T}}\;\eqqcolon\;1-p_{T}.$ (6)
On this event, for all $t\geq m_{T}$ we have $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by Eq. (5),
$\big|V_{(N)}(0^{+})-V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\big|<\varepsilon.$
For the first $m_{T}$ rounds we have $|V_{(N)}(0^{+})-V_{(N)}(\cdot)|\leq B$ due to bounded value. Therefore, conditioning on the event in (6),
$\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon.$
Taking expectations and using that the per-round gap is $\leq B$ on the complement of the event in (6), we have
$\mathbb{E}\bigg[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\bigg]\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon+B\,p_{T}$
$\;\leq\;\frac{O(\sqrt{T}\log\tfrac{1}{\delta})}{T}\,B+\varepsilon+BNTe^{-cO(\sqrt{T}\log\tfrac{1}{\delta})}~\leq~3{\varepsilon}$
for sufficiently large $T$. So we obtain
$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]\,=\,V_{(N)}(0^{+}).$
Combining with Lemma 6, we conclude that
$V_{(N)}(0^{+})\;\geq\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{I}(\lambda_{1,t}^{I},\ldots,\lambda_{N,t}^{I})\Big]\;=\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{I}\Big],$
which proves Theorem 4.1. ∎
5 Extension to Non-i.i.d. Valuations
We relax the i.i.d. assumption and allow bidder-specific value distributions. In each round $t\in[T]$, bidder $i\in[N]$ draws a value $v_{i,t}\sim F_{i}$ (independently across bidders and rounds), and the outside bid $d^{O}_{t}$ is drawn from $D$. Values and outside bids are bounded by $B>0$. Let $V_{t}^{C}$ and $V_{t}^{I}$ denote the coalition’s total value achieved in round $t$ under coordinated and independent bidding, respectively. Let $U_{i}^{\cdot,A}$ denote the cumulative RoS violation (achieved value minus payment) of bidder $i$ over $T$ rounds under scenario ${\cdot}\in\{C,I\}$ when all bidders follow auto-bidding algorithm $A$.
Total RoS violation.
Our first result shows that, in the non-i.i.d. setting, coordination strictly reduces the total RoS constraint violation, namely, the total utility of the coalition bidders is improved:
Theorem 5.1(Coordination reduces Total RoS Violation, non-iid case)
Assume 1. For any auto-bidding algorithm $A$ that (weakly) overbids, with probability at least $1-\exp\!\big(-\tfrac{\Delta^{2}T}{32B^{2}}\big)$, the total utility $\sum_{i=1}^{N}U_{i}^{C,A}\geq\sum_{i=1}^{N}U_{i}^{I,A}+\frac{\Delta T}{2}$.
See Appendix 0.B for the proof.
Total value.
To compare total value, we require a per-bidder margin condition that ensures each bidder wins with a positive average advantage when they are the highest-value coalition member. For bidder $i$, define $\Delta_{i}:=\mathbb{E}\left[(v_{i}-d^{O})\,\mathbb{I}\{v_{i}=\max_{j\in[N]}v_{j}\}\right]$.
Assumption 2 (Positive per-bidder winning margin)
For every bidder $i\in[N]$, $\Delta_{i}>0$.
Theorem 5.2(Coordination Increases Total Value, non-iid case)
Assume 2, and suppose bidders run Algorithm 3. Then, we have $\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0$.
See Appendix 0.C for the proof.
The findings in this section suggest that coordination improves upon independent bidding even in the presence of asymmetry inside the coalition, demonstrating the robustness of our conclusions.
Figure Setting NN $T$ Total Utility (I) Total Utility (C) Total Value (I) Total Value (C)
1(a) i.i.d. 2 4000 -0.011 0.220 0.643 0.666
1(b) i.i.d. 4 4000 -0.077 0.302 0.773 0.800
1(c) i.i.d. 3 4000 -0.048 0.154 0.713 0.749
2(a) non-i.i.d. 2 10000 0.049 0.219 0.715 0.718
2(b) non-i.i.d. 3 20000 -0.014 0.258 0.620 0.633
2(c) non-i.i.d. 5 20000 -0.062 0.619 0.814 0.819
3(a) Real data 4 20000 -0.040 0.156 $\pm$ 0.012 0.614 $\pm$ 0.016 0.929
3(b) Real data 5 20000 -0.065 0.174 $\pm$ 0.012 0.607 $\pm$ 0.012 0.959
Table 1: Performance summary. Confidence intervals below $\pm 0.002$ are omitted.
(a) $N=2,F=U[0,1],D=U[0,0.9]$
(b) $N=4,F=U[0,1],D=U[0,1]$
(c) $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$
Figure 1: Experiments under i.i.d auto-bidders.
6 Experiments
To complement our theoretical analysis, we conduct a series of experiments evaluating the performance of coordinated auto-bidding, on both synthetic data and a public real-world dataset (iPinYou [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou]).
Unless otherwise stated, all metrics are averaged over 100 independent simulations; shaded regions in figures depict 95% confidence intervals computed as $\bar{x}\pm 1.96\cdot s$, where $\bar{x}$ and $s$ are the sample mean and standard deviation over the 100 simulations.
6.1 Experiment Setup
For all experiments, we run Algorithm 3 with learning rate $\alpha=1/\sqrt{T}$, where $T$ is the total number of auction rounds. The horizon $T$ may vary across experiments because different settings exhibit different convergence rates. The values of $T$ and $N$ (number of coalition bidders) used in each setting are summarized in Table 1.
6.1.1 Synthetic Data
Symmetric setting.
Each coalition bidder $i$ draws values i.i.d. from uniform distribution $U[0,1]$. The outside bid is sampled from $D=U[0,0.9]$ (Figure 1(a)), $D=U[0,1]$ (Figure 1(b)), or $D=\mathrm{Beta}(3,2)$ (Figure 1(c)), representing different market conditions faced by the coalition bidders.
Asymmetric setting.
To stress-test the mechanism under heterogeneity, we sample each bidder’s values from distinct families (Uniform, Beta, or truncated Gaussians with bidder-specific parameters), as summarized in Figure 2. The outside-bid distribution $D$ varies with the scenario: $U[0.2,0.8]$ in Figure 2(a), $\mathrm{Beta}(3,5)$ in Figure 2(b), and $\mathrm{Beta}(2,8)$ in Figure 2(c). Asymmetries arise from both inside the coalition and the outside market conditions.
Full distributional specifications for the synthetic experiments are provided in the figure captions (Figures 1 and 2). All distributions conform to Assumption 1.
6.1.2 Real-World Data
We also experimented with the public iPinYou dataset [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou], released for a real-time bidding competition organized by iPinYou. We use the Season 2 test split, which contains 2,521,6302{,}521{,}630 auction records from $5$ advertisers. Each record provides the bidding (winning) price and the corresponding advertiser ID.
The real-world data impose three limitations: (i) only the winning price is observed, (ii) bidder identities are unavailable, and (iii) there is no label indicating whether any bidders coordinate. To obtain bidder valuations consistent with the data while abstracting away unobserved identities and strategies, we adopt an empirical i.i.d. model: we aggregate all observed winning prices into a common pool and, in each round, draw each coalition bidder’s value independently and uniformly at random with replacement from this pool. This is equivalent to sampling i.i.d. values from the empirical price distribution induced by the dataset.
Following prior work [chen_coordinated_2023], we normalize all prices to $[0,1]$ for numerical stability. The outside bidder’s bid is drawn from the same empirical distribution and then scaled by a random factor from $U[1,2]$ to maintain competitive pressure. We report results for coalitions with $N\in\{4,5\}$ bidders under this setup.
6.2 Results
Metrics.
Across all experiments, we evaluate three metrics for both Independent (Algorithm 1) and Coordination (Algorithm 2): (a) utility (achieved value minus payment), (b) achieved value, and (c) multiplier. Higher utility indicates lower RoS violation; in particular, utility $\geq 0$ implies that the RoS constraint is satisfied. Figures report per-bidder trajectories, whereas Table 1 aggregates to coalition-level totals (sum across bidders). All statistics are averaged over 100 independent runs.
Plot notation.
In the figures, labels of the form bidder $i$ -- I denote bidder $i$ under the Independent scenario, whereas bidder $i$ -- C denotes the Coordination scenario. Shaded regions show 95% confidence intervals over 100 simulations.
Results on synthetic data
In the symmetric settings (Figure 1), coordination sharply increases both each bidder’s utility (hence reducing RoS violation) and the coalition’s total achieved value relative to independent bidding. Although our theory guarantees value improvement at the coalition (total) level, the experiments further show that every individual bidder benefits from coordination by obtaining higher value. Moreover, all multipliers converge to 0, consistent with our theory, and per-bidder statistics (violation, value, and multiplier trajectories) closely track one another due to the i.i.d. value draws.
In the asymmetric settings (Figure 2), the aggregate gains largely persist: coordination improves total value on average and typically improves violations. However, heterogeneity induces dispersion in learning dynamics – bidders with different value distributions exhibit different convergence rates of the multiplier, and individual value improvements are not guaranteed for every bidder even though the coalition-level total value increases. In cases where each bidder has a non-negligible chance to be the winning bidder, we also observe that the multipliers tend to converge toward 0 empirically.
Results on real-world data.
Consistent with the synthetic data experiments, real-world data results show improvements in both individual RoS violation and achieved value, accompanied by multipliers converging steadily toward 0.
7 Discussion
In this paper, we studied coordinated auto-bidding mechanisms for value maximization under RoS constraints, to understand how cooperation among learning agents enhances efficiency in repeated auctions. We proposed a simple yet effective coordination protocol and theoretically proved that coordination improves the coalition’s total value, as well as RoS constraint compliance, under mild assumptions. Our empirical results further validate these findings, showing higher achieved value and consistent convergence patterns. The demonstrated benefit of coordination over independent bidding might provide useful guidance for industrial practice.
Building on these findings, we outline several directions for future work:
  • Auto-bidding outside bidders: While we assumed i.i.d. outside bids (a reasonable assumption when outside bidders are uniformly sampled from a population every round), a natural extension is to incorporate auto-bidding outside bidders, allowing both coalition members and external participants to adjust their bids dynamically.
  • Alternative coordination mechanisms: Another direction is to explore alternative coordination mechanisms, such as highest-bid coordination, to examine how they would affect the bidders’ overall welfare. Highest-bid coordination is analogous to highest-value coordination in the symmetric setting, but not in the asymmetric setting.
  • Other auction formats: Finally, it would be valuable to extend our analysis to other auction formats, especially non-truthful auctions such as first-price and generalized second-price auctions. Such extensions would help test the robustness of our theoretical insights and clarify how coordination interacts with strategic bidding incentives in broader market environments.
Appendix 0.A Proof of Lemma 3
Write the single-bidder’s expected utility as $G(\lambda)=\mathbb{E}_{v\sim F_{(N)}}G(\lambda;v)$ where $G(\lambda;v)$ is the bidder’s expected utility conditioning on having value $v$, which is equal to
$G(\lambda;v)$ $=\mathbb{E}_{d^{O}\sim D}\big[(v-d^{O})\cdot\mathbb{I}[(1+\tfrac{1}{\lambda})v>d^{O}]\big]$
$=vx\big((1+\tfrac{1}{\lambda})v\big)-p\big((1+\tfrac{1}{\lambda})v\big),$
where $x(b)$ is the winning probability when the bidder bids $b$ and $p(b)$ is the expected payment when the bidder bids $b$. Since the second-price auction is a truthful auction, by Myerson’s payment identity for truthful auctions, we have
$p(b)=xb-\int_{0}^{b}x(z)\mathrm{d}z.$ (7)
So,
$G(\lambda;v)$
$=vx\big((1+\tfrac{1}{\lambda})v\big)-(1+\tfrac{1}{\lambda})vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz$
$=-\tfrac{1}{\lambda}vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz.$
Take derivative with respect to $\lambda$:
$\tfrac{\partial G(\lambda;v)}{\partial\lambda}$
$=\tfrac{1}{\lambda^{2}}vx\big((1+\tfrac{1}{\lambda})v\big)-\tfrac{1}{\lambda}vx^{\prime}\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})+x\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})$
$=\tfrac{v^{2}}{\lambda^{3}}x^{\prime}\big((1+\tfrac{1}{\lambda})v\big)$
$\geq 0$
where the last inequality is because $x^{\prime}(b)\geq 0$. Integrating over $v\sim F_{(N)}$, we obtain $G^{\prime}(\lambda)\geq 0$, hence $G(\lambda)$ is increasing.
Appendix 0.B Proof of Theorem 5.1
The argument parallels the proof of Theorem 3.1.
Let $i^{*}_{t}=\arg\max_{i\in[N]}v_{i,t}$ denote the index of the highest-value bidder in round $t$. Conditioned on the history $H_{t-1}$, we analyze the expected increment $\Delta_{i,t}=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$ for bidder $i$:
$\mathbb{E}_{\bm{v}_{t}\sim F,\,d_{t}^{O}\sim D}\!\left[\Delta_{i,t}\mid H_{t-1}\right]$
$=\mathbb{E}_{\bm{v}_{t},\,d_{t}^{O}}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\Delta_{i,t}\mid H_{t-1}\right]$
$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
$\geq\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\,\big(\max_{j\neq i}v_{j,t}-d_{t}^{O}\big)_{+}\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}]\cdot(v_{i,t}-d_{t}^{O})\right].$
Summing over all bidders $i\in[N]$, the indicator $\mathbb{I}[i^{*}_{t}=i]$ ensures that only the highest-value bidder contributes in each round. Hence,
$\mathbb{E}\Big[\sum_{i=1}^{N}\Delta_{i,t}\;\big|\;H_{t-1}\Big]$ $\geq\mathbb{E}\big[(v_{(N-1)}-d_{t}^{O})_{+}\big]-\mathbb{E}\big[(d_{t}^{O}-v_{(N)})_{+}\big]$
$\geq\Delta,$
where the second inequality follows from Assumption 1. Finally, applying Azuma’s inequality (in a similar way as in the proof of Theorem 3.1) completes the proof.
Appendix 0.C Proof of Theorem 5.2
The argument parallels the proof of Theorem 4.1.
Let $G_{i}^{C}(\lambda_{i})$ denote bidder $i$’s expected utility when $v_{i}\sim F_{i}$ and the coalition uses multiplier $\lambda_{i}>0$:
$G_{i}^{C}(\lambda_{i})\;=\;\mathbb{E}_{v_{1}\sim F_{1},\dots,v_{N}\sim F_{N},\;d^{O}\sim D}\big[(v_{i}-d^{O})\,x_{i}(\bm{v},\lambda_{i})\big],$
where $x_{i}(\bm{v},\lambda_{i})=1$ if $v_{i}=\max_{j\in[N]}\{v_{j}\}$ and $(1+\tfrac{1}{\lambda_{i}})v_{i}>d^{O}$, and 0 otherwise.
By reusing the proof of Lemma 3 (Appendix 0.A) with $D$ replaced by the distribution of $\max\{d^{O},\max_{j\neq i}v_{j}\}$, we claim that $G_{i}^{C}(\lambda_{i})$ is increasing in $\lambda_{i}$. Hence,
$G_{i}^{C}(\lambda_{i})\geq\lim_{\mu\to 0^{+}}G_{i}^{C}(\mu)=\mathbb{E}\!\left[(v_{i}-d^{O})\,\mathbb{I}\!\left\{v_{i}=\max_{j\in[N]}\{v_{j}\}\right\}\right]=:\Delta_{i},$
where the limit exists because $G_{i}^{C}(\cdot)$ is bounded and monotone.
We now present the parallel version of Lemma 5.
Lemma 7
Under Assumption 2, for any $i\in[N]$ and $t\geq 1$, with probability at least $1-\exp\!\big(-\tfrac{\Delta_{i}^{2}t}{32B^{2}}\big)$,
$\lambda_{i,t+1}^{C}\;\leq\;\lambda_{i,1}\,\exp\!\Big(-\frac{\Delta_{i}\,t}{2\sqrt{T}}\Big).$
Proof
The expected utility $\mathbb{E}[g_{i,t^{\prime}}]$ is lower bounded by
$\mathbb{E}[g_{i,t^{\prime}}]\;=\;G_{i}^{C}(\lambda_{i,t^{\prime}})\;\geq\;\Delta_{i}.$
Applying Azuma’s inequality (exactly as in the proof of Lemma 5) completes the argument. ∎
Write $v_{(N)}\coloneqq\max_{j\in[N]}v_{j}$ with $v_{j}\sim F_{j}$ mutually independent. Its CDF is $F^{\max}(x)=\Pr[v_{(N)}\leq x]=\prod_{j=1}^{N}F_{j}(x)$. For $\lambda>0$, define
$V_{(N)}(\lambda)$ $\;\coloneqq\;\mathbb{E}_{\,v_{(N)}\sim F^{\max},\,d^{O}\sim D}\Big[\,v_{(N)}\cdot\mathbb{I}\big\{(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big\}\Big],$
$V_{(N)}(0^{+})$ $\;\coloneqq\;\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda).$
The limit $V_{(N)}(0^{+})=\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$ exists because $V_{(N)}(\lambda)$ is monotone and bounded. In the i.i.d. special case $F_{1}=\cdots=F_{N}=F$, $F^{\max}$ coincides with the order-statistic distribution $F_{(N)}$, recovering the original definition.
In the non-i.i.d. case, Lemma 6 still holds:
Lemma 8(copy of Lemma 6)
$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$.
Proof(Proof of Theorem 5.2)
Fix $\varepsilon>0$ and choose $\delta=\delta(\varepsilon)>0$ as in Equation (5). Let
$\Delta_{\min}\coloneqq\min_{i\in[N]}\Delta_{i}>0,\qquad c_{\min}\coloneqq\frac{\Delta_{\min}^{2}}{32B^{2}}$
Define
$m_{T}\;\coloneqq\;\Big\lceil\frac{2\sqrt{T}}{\Delta_{\min}}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\Big\rceil.$
By Lemma 7, for each $i$ and every $t\geq m_{T}$, $\lambda_{i,t+1}^{C}\leq\delta$ holds with probability at least $1-\exp(-c_{\min}t)$. A union bound over $i=1,\dots,N$ and $t=m_{T},\dots,T$ yields
$\Pr\!\Big[\bigcap_{i=1}^{N}\bigcap_{t=m_{T}}^{T}\{\lambda_{i,t+1}^{C}\leq\delta\}\Big]\;\geq\;1-NT\,e^{-c_{\min}m_{T}}\;\eqqcolon\;1-p_{T}.$
On this event, for all $t\geq m_{T}$ we have $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by (5),
$\big|V_{(N)}(0^{+})-V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\big|<\varepsilon,$
where $i^{*}_{t}=\arg\max_{i\in[N]}v_{i,t}$. For each of the first $m_{T}$ rounds, the absolute gap is $\leq B$. Therefore,
$\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon$
on the above high-probability event; on its complement the gap is $\leq B$. Taking expectations and using $p_{T}\leq NTe^{-c_{\min}m_{T}}$ gives
$\mathbb{E}\Big[\Big|V_{(N)}(0^{+})-\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\Big]\;\leq\;\frac{m_{T}}{T}B+\varepsilon+B\,p_{T}\;\xrightarrow[T\to\infty]{}\;0.$
Hence
$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]=V_{(N)}(0^{+}).$
Finally, combining with Lemma 8 yields the desired inequality against the independent-bidding benchmark, completing the proof. ∎
Appendix 0.D Additional Experiment Results
This appendix presents additional figures in the experiments: Figure 2 for the asymmetric setting on synthetic data; Figure 3 for the real-world dataset.
(a) $F_{1}=\mathrm{Beta}(2,5),F_{2}=\mathrm{Beta}(5,2),D=U[0.2,0.8],N=2$
(b) $F_{1}=\mathrm{Beta}(2,6),F_{2}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.6,0.15),F_{3}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.4,0.2),D=\mathrm{Beta}(3,5),N=3$
(c) $F_{1}=U[0.2,0.8],F_{2}=\mathrm{Beta}(4,3),F_{3}=\mathrm{Beta}(6,2),F_{4}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.5,0.1),F_{5}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.7,0.12),D=\mathrm{Beta}(2,8),N=5$
Figure 2: Experiments under non-i.i.d auto-bidders.
(a) $N=4$
(b) $N=5$
Figure 3: Experiments in real-world datasets.

### v2

@@
-
+On the Coordination of Value-Maximizing Bidders
 Abstract
 While the auto-bidding literature predominantly considers independent bidding, we investigate the coordination problem among multiple auto-bidders in online advertising platforms.
-Two motivating scenarios are: collaborative bidding among multiple distinct bidders managed by a third-party bidding agent, and strategic bid selection for multiple ad campaigns managed by a single advertiser.
-We formalize this coordination problem as a theoretical model and demonstrate that a straightforward coordination mechanism, where only the highest-value bidder competes with outside bids, strictly dominates independent bidding, improving both Return-on-Spend (RoS) compliance and the total value accrued for each participating auto-bidder or ad campaign.
-Additionally, our simulations on synthetic and real-world datasets support the theoretical result that coordinated mechanism outperforms independent bidding.
-These findings highlight both the theoretical potential and the practical robustness of coordination in auto-bidding in online auctions.
+Two motivating scenarios are: collaborative bidding among multiple bidders managed by a third-party bidding agent, and strategic bid selection for multiple ad campaigns managed by a single advertiser.
+We formalize this coordination problem as a theoretical model and investigate the coordination mechanism where only the highest-value bidder competes with outside bidders, while other coordinated bidders refrain from competing.
+We demonstrate that such a coordination mechanism dominates independent bidding, improving both Return-on-Spend (RoS) compliance and the total value accrued for the participating auto-bidders or ad campaigns, for a broad class of auto-bidding algorithms.
+Additionally, our simulations on synthetic and real-world datasets support the theoretical result that coordination outperforms independent bidding.
+These findings highlight both the theoretical potential and the practical robustness of coordinated auto-bidding in online auctions.
 1 Introduction
-Online advertising has been demonstrated as a prominent approach to monetize search engines, news feeds, social media, and e-commerce.
-Among them, real-time auctions have been used widely to connect advertisers and users efficiently.
-In the modern online advertising platforms, advertisers endeavor to maximize their campaign effectiveness, quantified by conversions or other pertinent metrics, through precise bid management that considers targeted return-on-spend (RoS) [aggarwal2019autobidding], which is also well-known as auto-bidding problems.
-To this end, a wide spectrum of bidding methodologies has been formulated, leveraging concepts from optimization theory, online learning paradigms, and game-theoretic principles [zhao2018deep, aggarwal2019autobidding, lee2013real, babaioff2021non, golrezaei2021bidding, deng2021towards, balseiro2021landscape, gao2022bidding, stram2024mystique].
+Online advertising is a prominent approach to monetize search engines, news feeds, social media, and e-commerce.
+Among them, real-time auctions have been widely used to connect advertisers and users efficiently.
+In modern online advertising platforms, advertisers endeavor to maximize their campaign effectiveness, quantified by conversions or other pertinent metrics, through precise bid management that considers targeted Return-on-Spend (RoS) (Aggarwal et al., 2019), which is also well-known as auto-bidding problem.
+To this end, a wide spectrum of bidding methodologies has been formulated, leveraging concepts from optimization theory, online learning paradigms, and game-theoretic principles (Zhao et al., 2018, Aggarwal et al., 2019, Lee et al., 2013, Babaioff et al., 2021, Golrezaei et al., 2021, Deng et al., 2021, Balseiro et al., 2021, Gao et al., 2022, Stram et al., 2024).
 Despite the wealth of sophisticated research in online advertising, existing literature predominantly concentrates on optimizing the value or utility from the perspective of individual bidders.
-This focus, while valuable, often overlooks a crucial aspect of real-world advertising dynamics: the potential for coordinated bidding strategies.
+This focus, while valuable, often overlooks a crucial aspect of real-world advertising dynamics: the possibility of coordinated bidding strategies.
 In practice, it is common for multiple bidders to form coalitions, perhaps by authorizing a third-party advertising platform to manage their bids collectively.
 Alternatively, a single large advertiser, such as major e-commerce players like Amazon, Temu, or Shein, might control numerous distinct advertising campaigns.
-These entities often have a portfolio of advertisements they are willing to display to users, necessitating a coordinated approach across their various campaigns to maximize overall effectiveness, rather than optimizing each in isolation.
+These entities often have a portfolio of advertisements they are willing to display to users, necessitating a coordinated approach across their various campaigns to maximize overall effectiveness, rather than optimizing each campaign in isolation.
 Consequently, a deeper understanding of how auto-bidding algorithms can be coordinated across multiple allied bidders or campaigns, and what are the benefits of coordination over independent bidding, is becoming increasingly critical.
-Such understandings might be useful for industry practices and unlock further efficiencies in online advertising markets.
+Such insights might be useful for industry practices and unlock further efficiencies in online advertising markets.
 In this work, we investigate the coordination problem among multiple auto-bidding algorithms in repeated auctions.
 We formulate a theoretical model in which $N$ auto-bidders participate in $T$ rounds of second-price auctions.
+The values of these bidders are independent and identically distributed (i.i.d.), motivated by the practice that only similar bidders are chosen to compete in ad auctions.
 These bidders may form a coalition to determine their bids jointly, competing against bidders outside the coalition.
-Each auto-bidder is modeled as a value-maximizing agent subject to a Return on Spend (RoS) constraint (i.e., non-negative overall utility).
+Each auto-bidder is modeled as a value-maximizing agent subject to an RoS constraint (i.e., non-negative overall utility).
 We consider a simple form of coordination: only the bidder with the highest value inside the coalition submits a positive bid to compete against the outside bidders, while the other N1N-1 coalition members refrain from competing.
@@
 The main contributions of this paper are as follows.
-Theoretically, we prove that the highest-value-bidder-compete coordination mechanism mentioned above, although straightforward, enjoys a significant advantage over independent bidding: By coordination, the bidders in the coalition acquire higher values, as well as achieve lower RoS constraint violations, compared to independent bidding.
-Specifically, we prove the following two results: (1) As long as the coalition bidders overbid (a common feature for value-maximizing bidding algorithms), then no matter what auto-bidding algorithms they use, coordination always increases the total utility (value - payment) of each bidder in the coalition over $T$ rounds, compared to the independent bidding scenario.
-This means that the bidders’ RoS constraints are better satisfied.
-(2) If coalition bidders run a state-of-the-art value-maximizing algorithm (a mirror-descent algorithm in feng2023online), then the total value obtained by the coalition is higher under coordination than under independent bidding.
-Our theoretical results are proven under two mild assumptions.
-The first one (Assumption 1) is that the highest and second-highest values among the bidders inside the coalition should be competitive against the highest bid outside of the coalition.
-Without such an assumption, the coalition as a whole is too weak to compete against the outside bidders, and the problem becomes non-interesting.
-The second assumption is that the bidders in the coalition have independent and identically distributed (i.i.d.) valuations.
-This i.i.d.
-assumption is not only for theoretical simplicity but also reasonable in practice.
-For the two motivating scenarios mentioned above (multiple bidders forming a coalition and a single bidder controlling multiple advertising campaigns), the bidders/campaigns selected by the auctioneer to compete usually share similar features, justifying i.i.d.
-valuations.
-In addition, we show in Section 5 that our analysis can be extended to non-i.i.d.
-settings.
-Finally, (in Section 6) we demonstrate by experiments, on both synthetic and real-world datasets, that our coordinated mechanism consistently outperforms independent bidding, validating our theoretical results under realistic market conditions.
+Theoretically, we prove that the highest-value-bidder-compete coordination mechanism mentioned above, although straightforward, enjoys a significant advantage over independent bidding: By coordination, bidders in the coalition acquire higher values, as well as achieve lower RoS constraint violations, compared to independent bidding, for a broad class of auto-bidding algorithms.
+In particular, we identify a condition on the distributions of bidders’ values and outside bids – Assumption 1 – under which each coalition bidder’s total utility in $T$ periods is higher, hence the bidder’s RoS constraint is less violated, under coordination than under independent bidding (Theorem 3.1).
+This conclusion holds for any auto-bidding algorithm that overbids (a common feature for value-maximizing algorithms).
+Assumption 1 is necessary and sufficient: when it does not hold, there exists an overbidding auto-bidding algorithm for which coordination hurts the coalition bidders, reducing their utility (RoS constraint compliance).
+We then consider the total value of coalition bidders.
+We prove that, as long as bidders use mirror-descent algorithms – a broad class of state-of-the-art auto-bidding algorithms under RoS constraints – then the total value of the coalition bidders will be improved by coordination, regardless of whether Assumption 1 holds (Theorem 4.1).
+If Assumption 1 holds, then such coordinated mirror-descent algorithms not only dominate independent bidding but also achieve the best possible coalition value among all possible coordination mechanisms, as $T\to\infty$ (Theorem 4.2).
+Beyond the i.i.d.
+setting, we extend our results to cases where coalition bidders’ values are independent but not identically distributed.
+In this non-i.i.d.
+setting, we demonstrate that – under slightly stronger assumptions – the same simple coordination algorithm still yields better total value and total RoS violation compared to independent bidding.
+But in contrast to the i.i.d.
+case, we cannot guarantee that this algorithm improves the performance for every individual bidder.
+Designing a coordination mechanism that achieves superior individual results in the non-i.i.d.
+setting remains an interesting open problem.
+Finally, Section 6 demonstrates, by experiments on synthetic and real-world datasets, that our coordinated mechanism consistently outperforms independent bidding, validating our theoretical results under realistic market conditions.
 1.1 Related Work
-Bidder coordination, often referred to as cartels or collusion in economics, has been extensively studied in traditional auction theory [robinson1985collusion, hendricks1989collusion, marshall2007bidder].
-Prior research has studied collusive behavior in first-price auctions [lopomo2011bidder, pesendorfer2000study], second-price auctions [mailath1991collusion, graham1987collusive], collusion-proof mechanism design [che2009optimal], auction design under collusion [pavlov2008auction], and collusion detection [chotibhongs2012analysis].
-While these studies typically consider bidders forming collusions autonomously and reaching equilibrium outcomes in static auction settings [bergemann2016bayes, fu2025learning], we instead introduce a central planner that coordinates bidders to improve their total welfare in dynamic auction environments.
-Recently, coordination behavior has also been observed in online advertising auctions [decarolis_bid_2023], sparking growing interest in understanding and analyzing such phenomena.
-decarolis2020marketing studied the impact of coordination on the revenue and efficiency of GSP and VCG auctions.
-romano2022power investigated the computational challenges faced by a media agency coordinating bidders under those mechanisms.
-chen_coordinated_2023 analyzed coordinated online bidding in repeated second-price auctions with budget constraints.
+Bidder coordination, often referred to as cartels or collusion in economics, has been extensively studied in traditional auction theory (Robinson, 1985, Hendricks and Porter, 1989, Marshall and Marx, 2007).
+Prior research has studied collusive behavior in first-price auctions (Lopomo et al., 2011, Pesendorfer, 2000), second-price auctions (Mailath and Zemsky, 1991, Graham and Marshall, 1987), collusion-proof mechanism design (Che and Kim, 2009), auction design under collusion (Pavlov, 2008), and collusion detection (Chotibhongs and Arditi, 2012).
+While these studies typically consider bidders forming collusions autonomously and reaching equilibrium outcomes in static auction settings (Bergemann and Morris, 2016, Fu and Lin, 2025), we instead introduce a central planner that coordinates bidders to improve their total welfare in dynamic auction environments.
+Recently, coordination behavior has also been observed in online advertising auctions (Decarolis et al., 2023), sparking growing interest in understanding and analyzing such phenomena.
+Decarolis et al.
+(2020) studied the impact of coordination on the revenue and efficiency of GSP and VCG auctions.
+Romano et al.
+(2022) investigated the computational challenges faced by a media agency coordinating bidders under those mechanisms.
+Chen et al.
+(2023) analyzed coordinated online bidding in repeated second-price auctions with budget constraints.
 While those works focus on coordinating utility maximizers, we study value maximizers.
@@
 Our work is also related to the large literature on auto-bidding.
-For example, paes2024complex studied the dynamics of systems with multiple auto-bidders and revealed complex behaviors such as bi-stability, periodic orbits, and quasi-periodicity, while aggarwal2025multi analyzed optimal bidding strategies for advertisers across multiple platforms.
-For a comprehensive overview of the literature, we refer readers to the survey by aggarwal_auto-bidding_2024.
+For example, Paes Leme et al.
+(2024) studied the dynamics of systems with multiple auto-bidders and revealed complex behaviors such as bi-stability, periodic orbits, and quasi-periodicity, while Aggarwal et al.
+(2025) analyzed optimal bidding strategies for advertisers across multiple platforms.
+For a comprehensive overview of the literature, we refer readers to the survey by Aggarwal et al.
+(2024).
 Our study complements the predominant individual-optimization perspective of the auto-bidding literature.
@@
 We consider the scenario where $N$ auto-bidders (or bidders) participate in $T$ rounds of repeated second-price auctions.
-In each round $t\in[T]=\{1,\ldots,T\}$, one item (e.g., ad slot) is for sale, and each bidder $i\in[N]=\{1,\ldots,N\}$ has a value $v_{i,t}\in[0,B]$ for the item i.i.d.
-sampled from a continuous distribution $F$.
+In each round $t\in[T]=\{1,\ldots,T\}$, one item (e.g., ad slot) is for sale, and each bidder $i\in[N]=\{1,\ldots,N\}$ draws a value $v_{i,t}\in[0,B]$ for the item i.i.d.
+from a continuous distribution $F$ with full support on $[0,B]$.
 Let $\bm{v}_{t}=(v_{1,t},\ldots,v_{N,t})$ be the vector of values of all bidders at round $t$.
@@
   1. 1.
-    NN advertisers, each with an ad, instruct their auto-bidding algorithms to coordinate with each other before bidding.
+    NN advertisers, each with an ad, instruct their auto-bidding algorithms to coordinate with each other.
   2. 2.
     One advertiser owns $N$ ads. Instead of using $N$ auto-bidding algorithms to bid for the $N$ ads independently, the advertiser instructs the algorithms to coordinate.
-In either case, usually in practice, only similar advertisers/ads are selected to compete for one ad slot, which justifies the i.i.d.
+In either case, usually in practice, only similar advertisers/ads are selected to compete for an ad slot, which justifies the i.i.d.
 value assumption.
@@
 sampled from another distribution $D$ every round.
-Our model also captures second-price auctions with reservation price, because $d_{t}^{O}$ can be the maximum between the reservation price and the outside competing bid.
+Our model also captures second-price auctions with reservation price, because $d_{t}^{O}$ can be the maximum of the reservation price and the outside competing bid.
 We denote the competing bid faced by bidder $i$ at round $t$ as
 $d_{i,t}=\max\big\{d^{O}_{t},\max_{j\in[N]\setminus\{i\}}b_{j,t}\,\big\}.$ (1)
-Let $x_{i,t}:=\mathbb{I}\{b_{i,t}\geq d_{i,t}\}\in\{0,1\}$ indicate whether bidder $i$ wins the item (ignoring tie-breaking for now), so bidder $i$’s payment is $p_{i,t}:=x_{i,t}d_{i,t}$ and utility is $u_{i,t}:=x_{i,t}(v_{i,t}-d_{i,t})$ at round $t$.
+Let $x_{i,t}:=\mathbb{I}\{b_{i,t}\geq d_{i,t}\}\in\{0,1\}$ indicate whether bidder $i$ wins the item (ignoring tie-breaking for now), so bidder $i$’s payment is $p_{i,t}:=x_{i,t}d_{i,t}$ and utility is $u_{i,t}:=x_{i,t}v_{i,t}-p_{i,t}=x_{i,t}(v_{i,t}-d_{i,t})$ at round $t$.
 Value Maximization under RoS Constraint.
 Each auto-bidder aims to maximize its total value while adhering to the Return on Spend (RoS) constraint. The formal definition of the optimization problem for bidder $i$ is as follows:
-$\max_{\{b_{i,t}\}_{t=1}^{T}}$ $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}$ (bidder $i$’s value)
-s.t. $\sum_{t=1}^{T}u_{i,t}=\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\sum_{t=1}^{T}p_{i,t}\geq 0$ (RoS constraint)
+$\max_{\{b_{i,t}\}_{t=1}^{T}}$ $\sum\nolimits_{t=1}^{T}v_{i,t}\cdot x_{i,t}$ (bidder $i$’s value)
+s.t. $\sum\nolimits_{t=1}^{T}\Big(\underbrace{v_{i,t}\cdot x_{i,t}-p_{i,t}}_{u_{i,t}}\Big)\geq 0$ $\text{(RoS constraint)}.$
 The RoS constraint ensures that for every dollar spent, at least one dollar of value is generated; it is equivalent to requiring the bidder’s total utility to be non-negative.
-[^1] In fact, the general RoS constraint requires $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\tau\cdot\sum_{t=1}^{T}p_{i,t}\geq 0$, i.e., at least $\tau$ dollar of value will be generated for every dollar spend, however, it is without loss of generality to assume $\tau=1$ in the above optimization problem because we can rescale the bidders’ values by dividing $\tau$.
-Independent and Coordinated Auto-Bidding
+[^1] In fact, the general RoS constraint requires $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\tau\cdot\sum_{t=1}^{T}p_{i,t}\geq 0$, i.e., at least $\tau$ dollar of value will be generated for every dollar spent.
+However, it is without loss of generality to assume $\tau=1$ in the above optimization problem because we can rescale the bidders’ values by dividing by $\tau$.
+Independent and Coordinated Auto-Bidding.
 Each auto-bidder runs an algorithm $A$ that learns to bid from history, in order to maximize the bidder’s total value subject to RoS constraint.
-Formally, at each round $t$, bidder $i$’s algorithm $A(\cdot)$ takes the historical information $H_{i,t}$ available to bidder $i$ (for example, the past values, bids, and allocations of the bidder: $H_{i,t}=\{v_{i,t^{\prime}},b_{i,t^{\prime}},x_{i,t^{\prime}}\}_{t^{\prime}=1}^{t-1}\cup\{v_{i,t}\}$) as input and outputs the bid $b_{i,t}$ for the current round.
-We refer to the scenario where the auto-bidding algorithms are run independently as independent bidding, formalized in Algorithm 1.
+Formally, at each round $t$, bidder $i$’s algorithm $A(\cdot)$ takes the historical information $H_{i,t}$ available to bidder $i$ – for example, the past values, bids, and allocations of the bidder: $H_{i,t}=\{v_{i,t^{\prime}},b_{i,t^{\prime}},x_{i,t^{\prime}}\}_{t^{\prime}=1}^{t-1}\cup\{v_{i,t}\}$ – as input and outputs the bid $b_{i,t}$ for the current round.
+We refer to the scenario where the auto-bidding algorithms are run independently as independent bidding, formalized below:
 for $t=1,2,\dots,T$ do
-   Each bidder $i\in[N]$ observes value $v_{i,t}$ and bids $b_{i,t}=A(H_{i,t})$;
-   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
+   Each bidder $i\in[N]$ observes their value $v_{i,t}\sim F$ and bids $b_{i,t}=A(H_{i,t})$.
+   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$.
 end for
@@
 We assume that ties are resolved naturally.
-[^2] Ties happen with probability 0 given a continuous value distribution $F$.
+[^2] Ties happen with probability 0 given a continuous $F$.
 This is formalized in Algorithm 2.
 for $t=1,2,\dots,T$ do
-   Each bidder $i\in[N]$ observes value $v_{i,t}$;
-   Let $i^{*}=\arg\max_{i\in[N]}v_{i,t}$;
-   Bidder $i^{*}$ bids $b_{i^{*},t}=A(H_{i^{*},t})$; all $i\neq i^{*}$ bid $b_{i,t}=0$;
-   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
+  Each bidder $i\in[N]$ observes their value $v_{i,t}\sim F$.
+  Let $i^{*}=\arg\max_{i\in[N]}v_{i,t}$.
+  Bidder $i^{*}$ bids $b_{i^{*},t}=A(H_{i^{*},t})$.
+  Bidder $i\neq i^{*}$ bids $b_{i,t}=0$.
+  Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$.
 end for
@@
 In this section, we analyze the extent to which the RoS constraint of every auto-bidder is maintained/violated.
-We will show that, under mild assumptions, coordinated bidding results in a smaller RoS violation compared to independent bidding.
-Equivalently, every bidder in the coalition obtains a higher utility through coordination.
-Formally, let $U_{i}^{C,A},U_{i}^{I,A}$ be the total utility of bidder $i$ during the $T$ rounds using auto-bidding algorithm $A$ under coordinated and independent bidding scenarios, respectively:
-$U_{i}^{C,A}=\sum_{t=1}^{T}u_{i,t}^{C,A},\quad U_{i}^{I,A}=\sum_{t=1}^{T}u_{i,t}^{I,A}.$ (2)
-RoS constraint requires $U_{i}^{C,A}\geq 0$ and $U_{i}^{I,A}\geq 0$. If $U_{i}^{C,A}\geq-c$ or $U_{i}^{I,A}\geq-c$, we say the RoS constraint for bidder $i$ is violated by $c$.
-We allow the bidders to use any auto-bidding algorithm that (weakly) overbids: namely, $A(\cdot)$ can be an arbitrary function that satisfies $b_{i,t}=A(H_{i,t})\geq v_{i,t}$.
-Note that we are considering value-maximizing bidders in second-price auctions.
-If the bidders underbid ($b_{i,t}<v_{i,t}$), they will always obtain non-negative utility and satisfy the RoS constraint, but not necessarily maximize their total values.
-So we assume the bidders to overbid.
-We make a reasonable assumption on the distributions of value and outside bid, $F$ and $D$.
+We will identify the necessary and sufficient condition under which coordinated bidding results in a smaller RoS violation compared to independent bidding (equivalently, every bidder in the coalition obtains a higher utility through coordination), for any auto-bidding algorithm.
+Formally, let $U_{i,T}^{C,A},U_{i,T}^{I,A}$ be the total utility of bidder $i$ during the $T$ rounds using auto-bidding algorithm $A$ under coordinated and independent bidding scenarios, respectively:
+$U_{i,T}^{C,A}=\sum\nolimits_{t=1}^{T}u_{i,t}^{C,A}\qquad U_{i,T}^{I,A}=\sum\nolimits_{t=1}^{T}u_{i,t}^{I,A}.$
+RoS constraint requires $U_{i,T}^{C\text{ or }I,A}\geq 0$. If $U_{i,T}^{C\text{ or }I,A}\geq-c$, we say the RoS constraint for bidder $i$ is violated by $c$.
+We allow the bidders to use any auto-bidding algorithm that (weakly) overbids: namely, $A(\cdot)$ can be any function that outputs bid $b_{i,t}=A(H_{i,t})\geq v_{i,t}$.
+Recall that we consider value maximization in second-price auctions.
+If bidders underbid ($b_{i,t}<v_{i,t}$), they will always obtain non-negative utility and satisfy the RoS constraint, but not necessarily maximize their values.
+So we assume overbidding.
+We introduce a condition/assumption on the value and outside bid distributions, $F$ and $D$.
 Let $(x)_{+}=\max\{x,0\}$ denote the positive part of a number.
@@
 samples from $F$.
-We assume that: in expectation, the advantage of the second-largest value $v_{(N-1)}$ over the outside bid $d^{O}$ is larger than the advantage of $d^{O}$ over the largest value $v_{(N)}$:
-Assumption 1
-$\mathbb{E}_{F,D}\big[\big(v_{(N-1)}-d^{O}\big)_{+}-\big(d^{O}-v_{(N)}\big)_{+}\big]\geq\Delta>0$.
-Intuitively, Assumption 1 says that the largest two values among the coalition bidders should be competitive enough against the outside bid.
-With large $v_{(N-1)}$ and $v_{(N)}$, $\mathbb{E}[(v_{(N-1)}-d^{O})_{+}]$ is large while $\mathbb{E}[(d^{O}-v_{(N)})_{+}]$ is small, which satisfy the assumption.
-Example 1
-The following examples satisfy Assumption 1:
+Assumption 1.
+In expectation, the advantage of the second-largest value $v_{(N-1)}$ over the outside bid $d^{O}$ is larger than the advantage of $d^{O}$ over the largest value $v_{(N)}$:
+$\Delta:=\mathbb{E}_{F,D}\big[(v_{(N-1)}-d^{O})_{+}-(d^{O}-v_{(N)})_{+}\big]\geq 0.$
+Intuitively, Assumption 1 says that the largest two values among the coalition bidders are competitive enough against the outside bid.
+This assumption is satisfied when the number $N$ of coalition bidders is large enough:
+Observation 1.
+For any full-support distributions $F$ and $D$, Assumption 1 is satisfied when $N$ is large enough.
+Assumption 1 is also satisfied by some natural distributions $F$ and $D$ with small $N$:
+Example 3.1.
+The following examples satisfy Assump. 1:
   • $N=4,F=U[0,1],D=U[0,1]$, with $\Delta=1/6$.
   • $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$, with $\Delta=1/40$.
-We will perform simulations on these examples in Section 6.
-The main result of this section is the following theorem, which shows that coordinated bidding strictly improves each bidder’s total utility compared to independent bidding.
-This directly implies that each bidder’s RoS constraint violation under coordination is smaller.
-Theorem 3.1(Coordination reduces RoS violation)
-Assume Assumption 1. For any auto-bidding algorithm $A$ that overbids, for every bidder $i\in[N]$, with probability at least $1-\exp(-\frac{\Delta^{2}T}{32B^{2}N^{2}})$,
-$U_{i}^{C,A}~\geq~U_{i}^{I,A}+\tfrac{\Delta T}{2N}.$
-The proof of Theorem 3.1 involves an interesting comparison between the coordinated bidding scenario and the truthful bidding scenario where all bidders bid their values exactly, and a comparison between truthful bidding and independent auto-bidding.
-These two comparisons together imply that coordinated bidding generates higher utility for every coalition member than independent bidding.
-We present the detailed proof below.
+The main result of this section is the following Theorem 3.1, which shows that Assumption 1 is the necessary and sufficient condition for coordination to reduce RoS constraint violation compared to independent bidding.
+When $\Delta\geq 0$, coordination improves the utility of every bidder in the coalition, for any auto-bidding algorithm that weakly overbids.
+When $\Delta<0$, this guarantee fails: coordination can be strictly worse than independent bidding for some overbidding algorithm.
+Theorem 3.1.
+If Assumption 1 holds (i.e., $\Delta\geq 0$), then for any (weakly) overbidding auto-bidding algorithm AA and every bidder $i\in[N]$,
+$\mathbb{E}\!\left[U_{i,T}^{C,A}-U_{i,T}^{I,A}\right]\geq\tfrac{T\Delta}{N}\geq 0.$
+If Assumption 1 fails (i.e., $\Delta<0$), then there exists an overbidding algorithm AA such that, for every bidder $i\in[N]$, for sufficiently large $T$,
+$\mathbb{E}\!\left[U_{i,T}^{C,A}-U_{i,T}^{I,A}\right]<0.$
+Remark 3.1.
+We can also obtain high-probability results from Theorem 3.1.
+Because values and outside bids are bounded by $B$, we can use Azuma’s inequality to prove that: when $\Delta>0$, for any overbidding algorithm $A$,
+$U_{i,T}^{C,A}~\geq~U_{i,T}^{I,A}+\tfrac{T\Delta}{2N}$
+holds with probability at least $1-\exp\big(-\tfrac{T\Delta^{2}}{32B^{2}N^{2}}\big)$.
 3.1 Proof of Theorem 3.1
-Let $U_{i}^{\mathrm{Truth}}=\sum_{t=1}^{T}u_{i,t}^{\mathrm{Truth}}$ be the total utility of bidder $i$ when all bidders in the coalition bid their values truthfully: $b^{\mathrm{Truth}}_{i,t}=v_{i,t}$.
-We will prove Theorem 3.1 by proving $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$ and $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$.
+The proof of the second (negative) direction is in Appendix A.2. Below we prove the first (positive) direction.
+Let $U_{i,T}^{\mathrm{Truth}}=\sum_{t=1}^{T}u_{i,t}^{\mathrm{Truth}}$ be the total utility of bidder $i$ when all bidders in the coalition bid their values truthfully: $b^{\mathrm{Truth}}_{i,t}=v_{i,t}$.
+We will prove Theorem 3.1 by proving $\mathbb{E}[U_{i,T}^{C,A}]\geq\mathbb{E}[U_{i,T}^{\mathrm{Truth}}]+\frac{\Delta T}{N}$ and $U_{i,T}^{\mathrm{Truth}}\geq U_{i,T}^{I,A}$.
 The second inequality is formalized below:
-Lemma 1
-For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$.
-Proof
+Lemma 3.1.
+For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, $U_{i,T}^{\mathrm{Truth}}\geq U_{i,T}^{I,A}$.
+Proof.
 This lemma directly follows from the dominant-strategy-incentive-compatibility of second-price auction, which ensures that each bidder’s truthful-bidding utility $u^{\mathrm{Truth}}_{i,t}$ is weakly larger than the non-truthful bidding utility $u^{I,A}_{i,t}$ at every round.
 ∎
-We then prove $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$:
-Lemma 2
+We then prove $\mathbb{E}[U_{i,T}^{C,A}]\geq\mathbb{E}[U_{i,T}^{\mathrm{Truth}}]+\frac{T\Delta}{N}$:
+Lemma 3.2.
+Under Assumption 1, for any auto-bidding algorithm that overbids, for every bidder $i\in[N]$,
 […]
-Proof
+Proof.
+See Appendix A.3. ∎
+Lemmas 3.1 and 3.2 together prove Theorem 3.1.
+4 Coordination Increases Bidders’ Value
+In this section, we show that coordination can increase the total value of auto-bidders, compared to independent bidding.
+This holds for a broad class of auto-bidding algorithms and without any distributional assumptions.
+In particular, we consider any mirror-descent-based auto-bidding algorithm.
+They are state-of-the-art algorithms for value maximization under RoS constraint, with near-optimal $O(\sqrt{T})$ regret guarantees (Balseiro et al., 2023, Feng et al., 2023).
+The algorithm maintains a parameter $\lambda_{i,t}$ for each bidder (which corresponds to the Lagrange multiplier of the dual optimization problem), and the bidder overbids by $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$ at each round.
+After observing a utility feedback, the bidder adjusts the parameter $\lambda_{i,t}$ in the opposite direction of the utility (i.e., positive utility decreases $\lambda_{i,t}$).
+The full algorithm is given in Algorithm 3.
+Definition 4.1 (Legendre mirror map and Bregman divergence).
+A $C^{1}$ function $h:(0,\infty)\to\mathbb{R}$ is a Legendre mirror map if: (i) $h$ is strictly convex on $(0,\infty)$; (ii) $h^{\prime}$ is a bijection from $(0,\infty)$ onto $\mathbb{R}$ (equivalently, $h^{\prime}$ is strictly increasing with $\lim_{\lambda\downarrow 0}h^{\prime}(\lambda)=$-$\infty$ and $\lim_{\lambda\uparrow\infty}h^{\prime}(\lambda)=+\infty$).
+The associated Bregman divergence is
+$D_{h}(\lambda,\mu):=h(\lambda)-h(\mu)-h^{\prime}(\mu)(\lambda-\mu),\quad\forall\lambda,\mu>0.$
+ Initialize multiplier $\lambda_{i,1}=1$, learning rate $\alpha=1/\sqrt{T}$.
+for $t=1,2,\cdots,T$ do
+  Observe value $v_{i,t}$ and set bid $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$.
+  Obtain allocation $x_{i,t}$ and pay $p_{i,t}$, compute realized utility $g_{i,t}=v_{i,t}\cdot x_{i,t}-p_{i,t}$.
+  Update the multiplier:
+$\lambda_{i,t+1}=\arg\min_{\lambda>0}\Big\{\alpha\cdot g_{i,t}\cdot\lambda+D_{h}(\lambda,\lambda_{i,t})\Big\}.$
+end for
+ALGORITHM 3 Mirror-Descent RoS Auto-Bidder (MD-$h$)
+As an example, Feng et al. (2023) study the mirror-descent-based algorithm with the entropy mirror map $h(\lambda)=\lambda\log\lambda-\lambda$, in which case the mirror-descent update is
+$\lambda_{i,t+1}=\lambda_{i,t}\exp(-\alpha g_{i,t}).$ (2)
+Feng et al. (2023)’s algorithm guarantees that each bidder’s total RoS violation in $T$ rounds is at most $O(\sqrt{T}\log T)$ and total value is at most $O(\sqrt{T})$ away from the optimum.
+Let
+$V_{t}^{C,A}=\sum\nolimits_{i=1}^{N}v_{i,t}x^{C,A}_{i,t}\qquad V_{t}^{I,A}=\sum\nolimits_{i=1}^{N}v_{i,t}x^{I,A}_{i,t}$
+be the total values obtained by the $N$ bidders at round $t$ in the coordinated and independent cases, respectively.
+We show that, for any mirror-descent-based algorithm $A$, coordination always improves the total value of coalition bidders.
+Theorem 4.1.
+Suppose bidders run a mirror-descent algorithm A (Algorithm 3). As $T\to\infty$, the coalition bidders’ total value under coordinated bidding is weakly larger than that under independent bidding:
+$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{C,A}\;-\;\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{I,A}\Big]\geq 0.$
+Proof sketch..
+We analyze the mirror-descent dynamics in the derivative space $y_{i,t}:=h^{\prime}(\lambda_{i,t})\in(-\infty,+\infty)$, where we can prove that the mirror-descent update rule is
+$y_{i,t+1}=y_{i,t}-\alpha g_{i,t}.$ (3)
+Let $G(\lambda)$ be the expected utility of a single bidder whose value is distributed according to the highest value $v_{(N)}\sim F_{(N)}$ among $N$ bidders, and who bids using multiplier $\lambda>0$, competing against the outside bid $d^{O}\sim D$:
+$G(\lambda)=\mathbb{E}\Big[(v_{(N)}-d^{O})\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
+Lemma 4.1.
+$G(\lambda)$ is strictly increasing in $\lambda>0$.
+Proof.
+See Lemma B.1. ∎
+Let $\lambda_{\star}>0$ be the solution to $G(\lambda_{\star})=0$ (we deal with the case of no solution in the formal proof seperately), and let $y_{\star}:=h^{\prime}(\lambda_{\star})$.
+The proof proceeds by identifying $\lambda_{\star}$ as a stable reference point for the coordinated mirror-descent dynamics and proving that $\lambda_{i,t}$ converges to $\lambda_{\star}$ in the coordinated case.
+We first characterize the expected update direction under coordination. We define $H_{t-1}:=\bigcup_{i=1}^{N}H_{i,t-1}$ as the joint history up to round t1t-1.
+Claim 4.1.
+Under coordinated bidding, conditioning on the history $H_{t-1}$, the expected realized utility (which serves as the gradient direction for multiplier update) satisfies
+$\mathbb{E}[g_{i,t}\mid H_{t-1}]=\tfrac{1}{N}G(\lambda_{i,t}).$
+Since $G(\lambda)$ is increasing, we have $G(\lambda_{i,t})\gtrless 0$ if and only if $\lambda_{i,t}\gtrless\lambda_{\star}$, giving expected gradient direction $\mathbb{E}[g_{i,t}\mid H_{t-1}]\gtrless 0$, so $y_{i,t}$ is decreased/increased towards $y_{\star}$ according to Eq.
+(3).
+This property allows us to view $(y_{i,t}-y_{\star})^{2}$ as a potential function, which decreases on average whenever $\lambda_{i,t}$ deviates significantly from $\lambda_{\star}$.
+By a potential function argument, we obtain:
+Claim 4.2.
+For any $\delta>0$, the fraction of rounds $t\leq T$ for which the active bidder $i_{t}^{\star}=\arg\max_{i}v_{i,t}$ satisfies $|y_{i_{t}^{\star},t}-y_{\star}|>\delta$ vanishes in expectation as $T\to\infty$:
+$\lim_{T\to\infty}\tfrac{1}{T}\sum\nolimits_{t=1}^{T}\Pr\bigl(|y_{i_{t}^{\star}}-y_{\star}|\geq\delta\bigr)=0.$
+The above claim implies that, asymptotically, the $y_{i,t}$ of every bidder $i\in[N]$ converges to $y_{\star}$, and hence the multiplier $\lambda_{i,t}$ converges to $\lambda_{\star}$.
+As a result, the coalition bidders’ total value converges to the expected value obtained by a single bidder whose value is $v_{(N)}\sim F_{(N)}$ and who bids by multiplier $\lambda_{\star}$, denoted by $V_{(N)}(\lambda_{\star})$:
+Claim 4.3.
+The coordinated time-average total value converges to the single bidder value
+$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{C,A}\Big]~=~V_{(N)}(\lambda_{\star}).$
+We then upper bound the time-average total value under independent bidding.
+Claim 4.4.
+Under independent bidding,
+$\limsup_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{I,A}\Big]~\leq~V_{(N)}(\lambda_{\star}).$
+To see this, note that for every bidder $i$ and all histories,
+$\mathbb{E}[g_{i,t}\mid H_{t-1}]\;\leq\;\tfrac{1}{N}G(\lambda_{i,t}),$
+and $G(\lambda_{i,t})<0$ when $\lambda_{i,t}<\lambda_{\star}$, where $y_{i,t}$ will be updated upwards.
+Consequently, no bidder can spend a non-vanishing fraction of rounds with $\lambda_{i,t}<\lambda_{\star}$, implying that independent bidders’ multipliers will eventually satisfy $\lambda_{i,t}\geq\lambda_{\star}$.
+Then because $V_{(N)}(\lambda)$ is monotonicaly decreasing, we obtain the claimed inequality $\mathbb{E}[V_{t}^{I,A}]\leq V_{(N)}(\lambda_{\star})$.
+Combining Claim 4.4 with Claim 4.3 yields
+$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{C,A}\Big]~\geq~\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{I,A}\Big],$
+which proves the theorem. Details are in Appendix B.1. ∎
+Theorem 4.1 does not require any assumption on the distributions of values or outside bid, except for boundedness and continuity.
+We further show that, if Assumption 1 holds, then the coordinated mirror-descent algorithm is not only better than the independent mirror-descent algorithm but also better than any coordination mechanism.
+Definition 4.2 (Coordination Mechanism).
+A coordination mechanism $\mathcal{G}$ is any procedure that, at each round $t$, given all realized values $\bm{v}_{t}=(v_{1,t},\ldots,v_{N,t})$, selects bids $(b_{1,t},\ldots,b_{N,t})$ for the coalition bidders.
+The mechanism may be history-dependent, i.e., bids at round $t$ may depend arbitrarily on past observations.
+Theorem 4.2.
+Under Assumption 1, as $T\to\infty$, the coalition bidders’ total value under coordinated mirror-descent algorithm AA (Algorithm 3) is weakly larger than that under any coordination mechanism $\mathcal{G}$:
+$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{C,A}\;-\;\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V_{t}^{\mathcal{G}}\Big]\geq 0,$
+where $V_{t}^{\mathcal{G}}$ is the total value obtained by the $N$ bidders at round $t$ using mechanism $\mathcal{G}$.
+Proof sketch..
+To prove Theorem 4.2, we first reduce the $N$-bidder coordinated scenario to a single-bidder problem where the single bidder’s value is the maximum among the coalition members.
+Using this reduction, we prove an interesting fact: under Assumption 1, the multipliers $\lambda_{i,t}^{C}$ of coordinated bidders converge to 0 in the limit as $t\to\infty$, meaning that the coordinated bidders will eventually submit arbitrarily high bids.
+This implies that the coordinated coalition will obtain the highest possible value, among all coordination mechanisms.
+The full proof is presented in Appendix B.3.
+∎
+5 Extension to Non-i.i.d. Valuations
+This section discusses the extension where bidders have independent but non-identically distributed values.
+In each round $t\in[T]$, bidders $i\in[N]$ draw values $v_{i,t}$ from different $F_{i}$.
+The outside bid $d^{O}_{t}$ is still drawn from $D$.
+Total RoS.
+In the i.i.d.
+case, coordination reduces the RoS constraint violation for every bidder in the coalition, if and only if Assumption 1 holds.
+In the non-i.i.d.
+case, we show that under the same assumption, coordination reduces the total RoS constraint violation among the coalition bidders, namely, the total utility of the coalition is improved.
+Recall that $U_{i,T}^{C,A}=\sum\nolimits_{t=1}^{T}u_{i,t}^{C,A}$ and $U_{i,T}^{I,A}=\sum\nolimits_{t=1}^{T}u_{i,t}^{I,A}$ are the total utilities of bidder $i$ during $T$ rounds under coordinated and independent bidding scenarios.
+Theorem 5.1 (Coordination reduces total RoS violation, non-i.i.d. case).
+Under Assumption 1 (i.e., $\Delta\geq 0$), for any overbidding algorithm $A$, the total utility among the $N$ coalition bidders satisfies
+$\mathbb{E}\Big[\sum\nolimits_{i=1}^{N}U_{i,T}^{C,A}\Big]~\geq~\mathbb{E}\Big[\sum\nolimits_{i=1}^{N}U_{i,T}^{I,A}\Big]+\Delta T.$
+The proof of this theorem is in Appendix C.1.
+Although Theorem 5.1 does not guarantee RoS improvement for every bidder, the total RoS improvement is also desirable in practice because the coalition bidders may redistribute the extra utility among themselves.
+Total value.
+In the i.i.d.
+setting, we showed that coordinated mirror descent algorithms achieve higher total value than independent mirror descent algorithms without any assumption, and higher than any coordination mechanism under Assumption 1.
+The proof, which relied on the fact that each bidder has the highest value among the coalition with probability $1/N$, does not apply to the non-i.i.d.
+case because bidders now differ in how often they are the highest-value coalition member.
+We therefore impose a per-bidder condition ensuring that each bidder contributes positively whenever they are selected.
+Concretely, for bidder $i$, define
+$\Delta_{i}:=\mathbb{E}\big[(v_{i}-d^{O})\,\mathbb{I}\{v_{i}=\max_{j\in[N]}v_{j}\}\big].$
+Assumption 2 (Positive per-bidder winning margin).
+For each bidder $i\in[N]$, $\Delta_{i}>0$.
+Theorem 5.2 (Coordination increases total value, non-i.i.d. case).
+Under Assumption 2, coordinated mirror descent algorithms (Algorithm 3) achieve weakly higher total value than independent mirror descent algorithms and any other coordination mechanism $\mathcal{G}$:
+$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V^{C,A}_{t}-\tfrac{1}{T}\sum\nolimits_{t=1}^{T}V^{\mathcal{G}}_{t}\Big]\geq 0.$
+See Appendix C.2 for the proof.
+The findings in this section suggest that coordination improves upon independent bidding even in the presence of asymmetry inside the coalition, demonstrating the robustness of our conclusions.
+Table 1: Performance summary. Confidence intervals no greater than $\pm 0.0025$ are omitted. All totals are normalized by $T$.
+Figure Setting NN $T$ Total Utility (I) Total Utility (C) Total Value (I) Total Value (C)
+1(a) i.i.d. 2 4000 -0.011 0.220 0.643 0.666
+3(a) i.i.d. 4 4000 -0.077 0.302 0.774 0.800
+3(b) i.i.d. 3 4000 -0.049 0.153 0.712 0.748
+1(b) non-i.i.d. 2 10000 0.049 0.219 0.715 0.718
+4(a) non-i.i.d. 3 20000 -0.014 0.258 0.619 0.633
+4(b) non-i.i.d. 5 20000 -0.062 0.619 0.814 0.819
+2 Real data 4 20000 -0.040 0.155 $\pm$ 0.012 0.620 $\pm$ 0.016 0.928 $\pm$ 0.003
+5 Real data 5 20000 -0.065 0.172 $\pm$ 0.012 0.608 $\pm$ 0.012 0.958
+6 Experiments
+To complement our theoretical analysis, we conduct a series of experiments evaluating the performance of coordinated auto-bidding, on both synthetic data and a public real-world dataset (iPinYou (Zhang et al., 2014, Liao et al., 2014)).
+Unless otherwise stated, all metrics are averaged over 100 independent simulations; shaded regions in figures depict 95% confidence intervals computed as $\bar{x}\pm 1.96\cdot s$, where $\bar{x}$ and $s$ are the sample mean and sample standard deviation divided by $\sqrt{100}$.
+6.1 Experiment Setup
+In all experiments, we run Algorithm 3 with mirror map $h(\lambda)=\lambda\log\lambda-\lambda$ Feng et al.
+(2023) (i.e., Equation˜2) and learning rate $\alpha=1/\sqrt{T}$, where $T$ is the total number of auction rounds.
+The horizon $T$ may vary across experiments because different settings exhibit different convergence rates.
+The values of $T$ and $N$ (number of coalition bidders) used in each setting are summarized in Table 1.
+6.1.1 Synthetic Data
+Symmetric setting.
+Each coalition bidder $i$ draws values i.i.d.
+from uniform distribution $U[0,1]$.
+The outside bid is sampled from $D=U[0,0.9]$ (Figure 1(a)), $D=U[0,1]$ (Figure 3(a)), or $D=\mathrm{Beta}(3,2)$ (Figure 3(b)), representing different market conditions faced by the coalition bidders.
+Asymmetric setting.
+To test the coordination mechanism under heterogeneity, we sample each bidder’s values from distinct families (Uniform, Beta, or truncated Gaussians with bidder-specific parameters).
+The outside-bid distribution $D$ varies with the scenario: $U[0.2,0.8]$ in Figure 1(b), $\mathrm{Beta}(3,5)$ in Figure 4(a), and $\mathrm{Beta}(2,8)$ in Figure 4(b).
+Asymmetries arise from both inside the coalition and the outside market conditions.
+Full distributional specifications for the synthetic experiments are provided in the figure captions (Figures 1(a), 3(a), 3(b), 1(b), 4(a), 4(b)). All distributions conform to Assumption 1.
+6.1.2 Real-World Data
+We also experimented with the public iPinYou dataset (Zhang et al., 2014, Liao et al., 2014), released for a real-time bidding competition organized by iPinYou.
+We use the Season 2 test split, which contains 2,521,6302{,}521{,}630 auction records from $5$ advertisers.
+Each record provides the bidding (winning) price and the corresponding advertiser ID.
+The real-world data impose three limitations: (i) only the winning price is observed, (ii) bidder identities are unavailable, and (iii) there is no label indicating whether any bidders coordinate.
+To obtain bidder valuations consistent with the data while abstracting away unobserved identities and strategies, we adopt an empirical i.i.d.
+model: we aggregate all observed winning prices into a common pool and, in each round, draw each coalition bidder’s value independently and uniformly at random with replacement from this pool.
+This is equivalent to sampling i.i.d.
+values from the empirical price distribution induced by the dataset.
+Constructing a reliable non-i.i.d.
+model would require persistent bidder identities or side information about coordination patterns, which are not available in the data.
+We normalize all prices to $[0,1]$ for numerical stability.
+The outside bidder’s bid is drawn from the same empirical distribution and then scaled by a random factor from $U[1,2]$ to maintain competitive pressure.
+We report results for coalitions with $N\in\{4,5\}$ bidders under this setup.
+6.2 Results
+Metrics.
+Across all experiments, we evaluate two metrics for both Independent (Algorithm 1) and Coordinated (Algorithm 2) algorithms: (a) utility (achieved value minus payment), (b) achieved value.
+Higher utility indicates lower RoS violation; in particular, utility $\geq 0$ implies that the RoS constraint is satisfied.
+Figures report per-bidder trajectories, whereas Table 1 aggregates to coalition-level totals (sum across bidders).
+All statistics are averaged over 100 independent runs.
+Plot notation.
+In the figures, labels of the form bidder $i$ - I denote bidder $i$ under the Independent scenario, whereas bidder $i$ - C denotes the Coordinated scenario.
+Shaded regions show 95% confidence intervals over 100 simulations.
+Results on synthetic data
+In the symmetric settings (Figures 1(a), 3(a), 3(b)), coordination sharply increases both each bidder’s utility (hence reducing RoS violation) and the coalition’s total achieved value relative to independent bidding.
+Although our theory only guarantees value improvement at the coalition (total) level, the experiments further show that every individual bidder benefits from coordination by obtaining higher value.
+We observe that per-bidder curves closely track one another due to the i.i.d.
+value draws.
+In the asymmetric settings (Figures 1(b), 4(a), 4(b)), the aggregate gains largely persist: coordination improves total value on average and typically improves violations.
+However, heterogeneity induces dispersion in learning dynamics – bidders with different value distributions exhibit different convergence rates, and individual value improvements are not guaranteed for every bidder even though the coalition-level total value increases.
+(a) Experiments under i.i.d. auto-bidders ($N=2,F=U[0,1],D=U[0,0.9]$).
+(b) Experiments under non-i.i.d. auto-bidders ($N=2,F_{1}=\mathrm{Beta}(2,5),F_{2}=\mathrm{Beta}(5,2),D=U[0.2,0.8]$).
+Figure 1: Experiments on synthetic data.
+Results on real-world data.
+Consistent with the synthetic data experiments, real-world data results (Figures 2, 5) show improvements in both individual RoS violation and achieved value.
+Figure 2: Experiments on real-world data. ($N=4$)
+7 Discussion
+In this paper, we studied coordinated auto-bidding mechanisms for value maximization under RoS constraints, to understand how cooperation among learning agents enhances efficiency in repeated auctions.
+We proposed a simple yet effective coordination protocol, theoretically identified the exact condition under which coordination improves the RoS compliance for every bidder, and proved that coordination unconditionally increases the total value attained by the coalition compared to independent bidding.
+Our empirical results further validate these findings, showing higher achieved value and lower RoS violation.
+The demonstrated benefits of coordination over independent bidding might provide useful guidance for industrial practice.
+Building on these findings, we outline several directions for future work:
+  • Auto-bidding outside bidders: While we assumed i.i.d.
+outside bids (a reasonable assumption when outside bidders are uniformly sampled from a population every round), a natural extension is to incorporate auto-bidding outside bidders, allowing both coalition members and external participants to adjust bids dynamically.
+  • Alternative coordination mechanisms: While we proved that our highest-value-compete coodination mechanism is asymptotically optimal under Assumption 1, it is still interesting to examine whether alternative mechanisms are better when the assumption does not hold.
+Moreover, if bidders’ private values are not accessible, then one may consider other coordination mechanisms, such as highest-bid-compete or mechanisms involving private information elicitation.
+  • Other auction formats: Finally, it would be valuable to extend our analysis to other auction formats, especially non-truthful auctions such as first-price and generalized second-price auctions.
+Such extensions would help test the robustness of our theoretical insights and clarify how coordination interacts with strategic bidding incentives in broader market environments.
+Appendix A Omitted Proofs in Section 3
+A.1 Proof of Observation 1
+Fix $\varepsilon\in(0,B)$.
+The probability $q_{\varepsilon}:=\Pr_{v\sim F}[v>B-{\varepsilon}]=1-F(B-\varepsilon)>0$ by full support/continuity of $F$.
+Let $M_{N}:=\max\{v_{1},\dots,v_{N}\}=v^{(N)}$ and $S_{N}:=v^{(N-1)}$.
+The events $\{v_{n}>B-\varepsilon\}$ are i.i.d.
+with probability $q_{\varepsilon}$ and $\sum_{n=1}^{\infty}\Pr(v_{n}>B-\varepsilon)=\infty$, so by the second Borel–Cantelli lemma, $v_{n}>B-\varepsilon$ occurs infinitely often for all sufficiently large $n$ with probability one.
+By monotonicity of the running maxima and of the second maximum, this implies $M_{N}\geq B-\varepsilon$ and $S_{N}\geq B-\varepsilon$ for all sufficiently large $N$ with probability one; hence $M_{N}\to B$ and $S_{N}\to B$ almost surely.
+Define $Z_{N}:=(S_{N}-d^{O})_{+}-(d^{O}-M_{N})_{+}\in[-B,B]$ and $\Delta_{N}=\mathbb{E}[Z_{N}]$.
+Since $M_{N},S_{N}\to B$ almost surely and $d^{O}\in[0,B]$, we have $Z_{N}\to(B-d^{O})_{+}$ almost surely.
+By the dominated convergence theorem,
+$\lim_{N\to\infty}\Delta_{N}~=~\lim_{N\to\infty}\mathbb{E}[Z_{N}]~=~\mathbb{E}[\lim_{N\to\infty}Z_{N}]~=~\mathbb{E}\!\left[(B-d^{O})_{+}\right]~\geq~0,$
+with strict inequality whenever $\Pr(d^{O}<B)>0$. Therefore there exists $N_{0}$ such that $\Delta_{N}\geq 0$ for all $N\geq N_{0}$.
+A.2 Proof of the Second (Negative) Result in Theorem 3.1
+Suppose Assumption 1 does not hold.
+Let the upper bound on value and outside bid be $B=1$.
+For every bidder $i\in[N]$, we construct the following (weakly) overbidding algorithm $A$ that chooses the current bid based on the bidder’s past bids:
+$A(H_{i,t})\;=\;\begin{cases}1,&\text{if there exists some }s<t\text{ with }b_{i,s}=0,\\ v_{i,t},&\text{otherwise.}\end{cases}$
+Claim A.1.
+For every horizon $T\geq 1$, every bidder $i$’s expected total utility difference is
+$\mathbb{E}\!\left[U_{i,T}^{C,A}-U_{i,T}^{I,A}\right]\;=\;\frac{T}{N}\,\Delta\;+\;\frac{1-N^{-T}}{N-1}\,L.$ (4)
+where $L\;:=\;\mathbb{E}\big[(d^{O}-v^{(N)})_{+}\big].$
+In particular, if $\Delta<0$ (i.e., Assumption 1 fails with strict negativity), then for all
+$T\;>\;\frac{N}{N-1}\cdot\frac{L}{-\Delta},$
+we have $\mathbb{E}[U_{i,T}^{C,A}]<\mathbb{E}[U_{i,T}^{I,A}]$.
+Proof.
+We first note that, under independent bidding, the first case of algorithm $A$ never triggers.
+Indeed, since $v_{i,t}>0$ almost surely (by continuity of $F$) and $A$ outputs either $v_{i,t}$ or $1$, bidder $i$ never submits a zero bid.
+Hence the condition $\exists s<t$ with $b_{i,s}=0$ is never satisfied, so $b_{i,t}=v_{i,t}$ for every $t\in[T]$ and every bidder $i\in[N]$, hence
+$U_{i,T}^{I,A}=U_{i,T}^{\mathrm{Truth}},$
+where $U_{i,T}^{\mathrm{Truth}}$ is bidder $i$’s total utility when all bidders in $[N]$ bid truthfully.
+We now analyze the coordinated scenario. In each round $t$, only the highest-value bidder $i_{t}^{*}$ follows $A$, while all other bidders bid 0. Fix a bidder $i$ and define
+$\tau_{i}:=\min\{t\geq 1:i\neq i_{t}^{*}\},$
+as the first time when bidder $i$ does not have the highest value, so that bidder $i$ bids 0 for the first time in round $\tau_{i}$, and the trigger is active from round $\tau_{i}+1$ onward.
+By symmetry of values and independence across rounds,
+$\Pr[\tau_{i}\geq t]=\Pr[i_{1}^{*}=i,\ldots,i_{t-1}^{*}=i]=\Big(\frac{1}{N}\Big)^{t-1}.$
+Let $\Delta_{i,t}:=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$. Since $U_{i,T}^{C,A}-U_{i,T}^{I,A}=\sum_{t=1}^{T}\Delta_{i,t}$, it suffices to compute $\mathbb{E}[\Delta_{i,t}]$.
+- If $\tau_{i}<t$, bidder $i$ has already bid 0 in the past, so when selected (i.e., having the highest value $v_{t}^{(N)}$) in round $t$ it bids $1$.
+A direct comparison with truthful bidding shows that on this event
+[…]
+$~=~\big(v_{t}^{(N)}-d_{t}^{O}\big)~-~\big(v_{t}^{(N)}-\max\big\{v_{t}^{(N-1)},d_{t}^{O}\big\}\big)\cdot\mathbb{I}\big[v_{t}^{(N)}>d_{t}^{O}\big]$
+$~=~\mathbb{I}\big[v_{t}^{(N)}>d_{t}^{O}\big]\cdot\big(\max\big\{v_{t}^{(N-1)},d_{t}^{O}\big\}-d_{t}^{O}\big)~+~\mathbb{I}\big[v_{t}^{(N)}\leq d_{t}^{O}\big]\cdot\big(v_{t}^{(N)}-d_{t}^{O}\big)$
+$~=~(v_{t}^{(N-1)}-d_{t}^{O})_{+}-(d_{t}^{O}-v_{t}^{(N)})_{+}.$
+By symmetry, bidder $i$ is selected with probability $1/N$, so
+$\mathbb{E}[\Delta_{i,t}\mid\tau_{i}<t]~=~\frac{1}{N}\Delta.$
+- If instead $\tau_{i}\geq t$, bidder $i$ has not yet bid 0, and therefore bids truthfully when selected.
+In this case the only difference relative to truthful bidding is the reduced payment under coordination, yielding $\Delta_{i,t}=(v_{t}^{(N-1)}-d_{t}^{O})_{+}$,
+$\mathbb{E}[\Delta_{i,t}\mid\tau_{i}\geq t]~=~\frac{1}{N}\mathbb{E}[(v^{(N-1)}-d^{O})_{+}]~=~\frac{\Delta+L}{N}.$
+Combining the two cases,
+$\mathbb{E}[\Delta_{i,t}]~=~\Pr[\tau_{i}<t]\frac{\Delta}{N}+\Pr[\tau_{i}\geq t]\frac{\Delta+L}{N}~=~\frac{\Delta}{N}+\frac{L}{N^{t}}.$
+Summing over $t=1,\ldots,T$ gives
+$\mathbb{E}[U_{i,T}^{C,A}-U_{i,T}^{I,A}]~=~\sum_{t=1}^{T}\mathbb{E}[\Delta_{i,t}]~=~\frac{T}{N}\Delta+\frac{1-N^{-T}}{N-1}L,$
+which proves Equation˜4. If $\Delta<0$, the final inequality follows immediately for $T$ exceeding the stated threshold. ∎
+A.3 Proof of Lemma 3.2
 For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, consider the difference between bidder $i$’s coordinated auto-bidding utility and truthful bidding utility at round $t$:
 $\Delta_{i,t}$ $=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$
-$=(v_{i,t}-d_{t}^{O})\cdot\mathbb{I}[b_{i,t}^{C,A}\geq d_{t}^{O}]\cdot\mathbb{I}[v_{i,t}=\max_{j\in[N]}v_{j,t}]$
-$\quad-(v_{i,t}-d_{i,t}^{\mathrm{Truth}})\cdot\mathbb{I}[v_{i,t}\geq d_{i,t}^{\mathrm{Truth}}]$
+$=(v_{i,t}-d_{t}^{O})\cdot\mathbb{I}[b_{i,t}^{C,A}\geq d_{t}^{O}]\cdot\mathbb{I}[v_{i,t}=\max_{j\in[N]}v_{j,t}]-(v_{i,t}-d_{i,t}^{\mathrm{Truth}})\cdot\mathbb{I}[v_{i,t}\geq d_{i,t}^{\mathrm{Truth}}]$
 where $d_{i,t}^{\mathrm{Truth}}=\max\{d_{t}^{O},\max_{j\neq i}v_{j,t}\}$ is the competing bid for bidder $i$ under truthful bidding. We do a case analysis:
@@
 Let $F_{(N)}$ be the distribution of the largest value $v_{(N)}$ among $N$ samples from $F$.
-[…]
-[…]
+$\mathbb{E}_{\bm{v}_{t}\sim F,d_{t}^{O}\sim D}\big[\Delta_{i,t}\mid H_{t-1}\big]$
+$=0+\Pr[v_{i,t}=\max_{j\in[N]}v_{j,t}]\cdot\mathbb{E}_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\big[\Delta_{i,t}\mid H_{t-1}\big]$
 $=\frac{1}{N}\bigg(\Pr_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\Big[v_{i,t}>d_{t}^{O}\Big]\cdot\mathbb{E}\Big[d_{t}^{\mathrm{Truth}}-d_{t}^{O}\,\big|\,v_{i,t}\geq d_{t}^{O}\Big]$
-$\quad+\Pr_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\Big[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\cdot\mathbb{E}\Big[v_{i,t}-d_{t}^{O}\,\big|\,v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\bigg)$
-$=\frac{1}{N}\bigg(\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}>d_{t}^{O}\big]\cdot\big(\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\big)\Big]$
-$\qquad+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg).$
+$\qquad+\Pr_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\Big[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\cdot\mathbb{E}\Big[v_{i,t}-d_{t}^{O}\,\big|\,v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\bigg)$
+$=\frac{1}{N}\bigg(\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}>d_{t}^{O}\big]\cdot\big(\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\big)\Big]+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
 Because $\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}>0$ implies $\mathbb{I}\big[v_{(N)}>d_{t}^{O}\big]=1$, the above is equal to
-$=\frac{1}{N}\bigg(\mathbb{E}\Big[\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\Big]$
-$\qquad+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
+$=\frac{1}{N}\bigg(\mathbb{E}\Big[\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\Big]+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
 $\geq\frac{1}{N}\bigg(\mathbb{E}\Big[\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\Big]+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
@@
 […]
-Because values and outside bids are bounded by $B$, $\Delta_{i,t}$ is bounded by [B,B][-B,B]. So, by Azuma’s inequality,
+Therefore,
+$\mathbb{E}\Big[U_{i,T}^{C,A}-U_{i,T}^{\mathrm{Truth}}\Big]~=~\mathbb{E}\left[\sum_{t=1}^{T}\Delta_{i,t}\right]~\geq~T\cdot\frac{\Delta}{N},$
+which proves the lemma.
+Appendix B Omitted Proofs in Section 4
+Denote the time-average total values among the $N$ coalition bidders by
+$V_{T}^{C,A}:=\frac{1}{T}\sum_{t=1}^{T}V_{t}^{C,A}\qquad V_{T}^{I,A}:=\frac{1}{T}\sum_{t=1}^{T}V_{t}^{I,A}$
+for the coordinated and independent cases, respectively.
+B.1 Proof of Theorem 4.1
+Let $v_{(N)}=\max\{v_{1},\ldots,v_{N}\}$ denote the highest value among $N$ i.i.d.
+samples from $F$.
+Let $F_{(N)}$ be the distribution of $v_{(N)}$.
+Let $G(\lambda)$ be the expected utility of a single bidder with value $v_{(N)}\sim F_{(N)}$ and bidding using multiplier $\lambda>0$, competing against the outside bid $d^{O}\sim D$:
+$G(\lambda)=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Big[(v_{(N)}-d^{O})\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
+We present the monotonicity property of $G(\lambda)$.
+Lemma B.1.
+$G(\lambda)$ is strictly increasing in $\lambda>0$.
+Proof.
+Write the single-bidder’s expected utility as $G(\lambda)=\mathbb{E}_{v\sim F_{(N)}}G(\lambda;v)$ where $G(\lambda;v)$ is the bidder’s expected utility conditioning on having value $v$, which is equal to
+$G(\lambda;v)~=~\mathbb{E}_{d^{O}\sim D}\big[(v-d^{O})\cdot\mathbb{I}\{(1+\tfrac{1}{\lambda})v>d^{O}\}\big]~=~v\cdot x\big((1+\tfrac{1}{\lambda})v\big)-p\big((1+\tfrac{1}{\lambda})v\big),$
+where $x(b)$ is the winning probability when the bidder bids $b$ and $p(b)$ is the expected payment when the bidder bids $b$. In particular,
+$x(b)=\int_{0}^{b}f_{D}(z)\mathrm{d}z,\quad p(b)=\int_{0}^{b}z\cdot f_{D}(z)\mathrm{d}z,$
+where $f_{D}(\cdot)$ is the density function of the outside bid distribution. Since the second-price auction is a truthful auction, by Myerson’s payment identity for truthful auctions, we have
+$p(b)=x(b)b-\int_{0}^{b}x(z)\mathrm{d}z.$ (5)
+So,
+$G(\lambda;v)$ $=vx\big((1+\tfrac{1}{\lambda})v\big)-(1+\tfrac{1}{\lambda})vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz$
+$=-\tfrac{1}{\lambda}vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz.$
+Take derivative with respect to $\lambda$:
+$\tfrac{\partial G(\lambda;v)}{\partial\lambda}$ $=\tfrac{1}{\lambda^{2}}vx\big((1+\tfrac{1}{\lambda})v\big)-\tfrac{1}{\lambda}vx^{\prime}\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})+x\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})$
+$=\tfrac{v^{2}}{\lambda^{3}}x^{\prime}\big((1+\tfrac{1}{\lambda})v\big)$
+$=\tfrac{v^{2}}{\lambda^{3}}f_{D}\big((1+\tfrac{1}{\lambda})v\big)$
+Integrating over $v\sim F_{(N)}$, we obtain
+$\tfrac{\mathrm{d}G(\lambda)}{\mathrm{d}\lambda}=\mathbb{E}_{v\sim F_{(N)}}\Big[\tfrac{v^{2}}{\lambda^{3}}f_{D}\big((1+\tfrac{1}{\lambda})v\big)\Big].$
+Because $F_{(N)}$ and $D$ are distributions with full support on $[0,B]$, we have $f_{D}(z)>0$ for $z\in[0,B]$ and the event $\tfrac{v^{2}}{\lambda^{3}}f_{D}\big((1+\tfrac{1}{\lambda})v\big)>0$ happens with positive probability.
+So, we have $\tfrac{\mathrm{d}G(\lambda)}{\mathrm{d}\lambda}>0$ and $G(\lambda)$ is strictly increasing.
+We now define the reference stable point
+$\lambda_{\star}\ :=\ \inf\big\{\lambda>0:\ G(\lambda)\geq 0\big\}\ \in\ [0,\infty).$ (6)
+We consider the generic case $\lambda_{\star}\in(0,\infty)$ here and discuss $\lambda_{\star}=0$ in Appendix B.2.
+Let $y_{i,t}:=h^{\prime}(\lambda_{i,t})$ and $y_{\star}:=h^{\prime}(\lambda_{\star})$.
+For the mirror-descent update
+$\lambda_{i,t+1}\in\arg\min_{\lambda>0}\big\{\alpha g_{i,t}\lambda+D_{h}(\lambda,\lambda_{i,t})\big\}$
+in Algorithm 3, first-order optimality yields $\alpha g_{i,t}+h^{\prime}(\lambda_{i,t+1})-h^{\prime}(\lambda_{i,t})=0$. Equivalently,
+$y_{i,t+1}=y_{i,t}-\alpha g_{i,t}.$ (7)
+Define $\phi(y):=\frac{1}{N}G\bigl((h^{\prime})^{-1}(y)\bigr)$.
+Since $(h^{\prime})^{-1}$ is strictly increasing and continuous and $G$ is strictly increasing (by Lemma B.1) and continuous, $\phi$ is continuous and strictly increasing.
+Step 1: coordination concentrates $y_{i,t}$ near $y_{\star}$.
+Under coordination, bidder $i$ is active iff $i=i_{t}^{*}$, which happens with probability $1/N$ since bidders are i.i.d. . Hence, conditioning on history $H_{t-1}$, bidder $i$’s expected utility is
+$\mathbb{E}[g_{i,t}\mid H_{t-1}]\;=\;\frac{1}{N}\,G(\lambda_{i,t})\;=\;\phi(y_{i,t}).$ (8)
+Fix $\delta>0$. Since $\phi$ is strictly increasing and $\phi(y_{\star})=0$, we have $\phi(y_{\star}-\delta)<0<\phi(y_{\star}+\delta)$. Define
+$c_{\delta}\ :=\ \min\bigl\{-\phi(y_{\star}-\delta),\ \phi(y_{\star}+\delta)\bigr\}\ >\ 0.$ (9)
+Then for all $y$ with $|y-y_{\star}|\geq\delta$,
+$(y-y_{\star})\,\phi(y)\ \geq\ \delta\,c_{\delta}.$ (10)
+Claim B.1.
+For every bidder $i\in[N]$ and any $\delta>0$,
+$\frac{1}{T}\sum_{t=1}^{T}\Pr\bigl(|y_{i,t}-y_{\star}|\geq\delta\bigr)\ \leq\ \frac{(y_{i,1}-y_{\star})^{2}+B^{2}}{2\delta c_{\delta}}\cdot\frac{1}{\sqrt{T}}.$ (11)
+Proof.
+Let $W_{t}:=(y_{i,t}-y_{\star})^{2}$. From Equation (7),
+$W_{t+1}=W_{t}-2\alpha(y_{i,t}-y_{\star})g_{i,t}+\alpha^{2}g_{i,t}^{2}.$
+Taking $\mathbb{E}[\cdot\mid H_{t-1}]$, using Equation (8) and $g_{i,t}^{2}\leq B^{2}$,
+$\mathbb{E}[W_{t+1}\mid H_{t-1}]\leq W_{t}-2\alpha(y_{i,t}-y_{\star})\phi(y_{i,t})+\alpha^{2}B^{2}.$
+Sum over $t=1,\dots,T$, use $W_{T+1}\geq 0$ and $\alpha^{2}T=1$:
+$2\alpha\sum_{t=1}^{T}\mathbb{E}\big[(y_{i,t}-y_{\star})\phi(y_{i,t})\big]\leq W_{1}+B^{2}.$
+By Equation (10), $\mathbb{E}\big[(y_{i,t}-y_{\star})\phi(y_{i,t})\big]\geq\delta c_{\delta}\cdot\Pr(|y_{i,t}-y_{\star}|\geq\delta)$. Plugging in and dividing by $2\alpha T$ gives Eq. (11). ∎
+Step 2: coordination value converges to $V_{(N)}(\lambda_{\star})$.
+Under coordination, the total value among the $N$ coalition bidders is
+$V_{T}^{C,A}=\frac{1}{T}\sum_{t=1}^{T}v_{(N),t}\,\mathbb{I}\left[\Bigl(1+\frac{1}{\lambda_{i_{t}^{\star},t}}\Bigr)v_{(N),t}>d_{t}^{O}\right].$
+Since $(v_{(N),t},d_{t}^{O})$ is independent of $H_{t-1}$,
+$\mathbb{E}[V_{T}^{C,A}]=\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i_{t}^{\star},t})\Big].$
+Fix $\varepsilon>0$.
+By continuity of $V_{(N)}$ at $\lambda_{\star}$, pick $\delta>0$ s.t.
+$|V_{(N)}(\lambda)-V_{(N)}(\lambda_{\star})|\leq\varepsilon$ whenever $|y_{i_{t}^{\star},t}-y_{\star}|<\delta$.
+Using $0\leq V_{(N)}(\cdot)\leq B$, we have
+$\left|\mathbb{E}[V_{T}^{C,A}]-V_{(N)}(\lambda_{\star})\right|\leq\varepsilon+\frac{2B}{T}\mathbb{E}\!\left[\sum_{t=1}^{T}\mathbb{I}\left[|y_{i_{t}^{\star},t}-y_{\star}|\geq\delta\right]\right]$ (12)
+By symmetry among the $N$ bidders, conditioning on history $H_{t-1}$,
+$\mathbb{E}\Big[\mathbb{I}\left[|y_{i_{t}^{*},t}-y_{\star}|\geq\delta\right]~\Big|~H_{t-1}\Big]=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[|y_{i,t}-y_{\star}|\geq\delta\right],$
+hence taking expectations and averaging over $t$,
+$\mathbb{E}\!\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left[|y_{i_{t}^{*},t}-y_{\star}|\geq\delta\right]\right]\leq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\!\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left[|y_{i,t}-y_{\star}|\geq\delta\right]\right]\to 0\quad\text{as}\quad T\to\infty$
+by Equation˜11. Therefore, first letting $T\to\infty$ in Equation˜12 and then $\varepsilon\downarrow 0$ gives
+$\lim_{T\to\infty}\mathbb{E}[V_{T}^{C,A}]\;=\;V_{(N)}(\lambda_{\star}).$ (13)
+Step 3: independent bidding has $\limsup$ value at most $V_{(N)}(\lambda_{\star})$.
+Fix bidder $i$ and round $t$ under independent bidding, and let $A_{i,t}$ be the event $\{i=i_{t}^{\star}\}$. For every realization, each bidder’s realized utility $g_{i,t}$ satisfies
+$g_{i,t}\ \leq\ \mathbb{I}\left[A_{i,t}\right]\,(v_{(N),t}-d_{t}^{O})\,\mathbb{I}\left[\Bigl(1+\frac{1}{\lambda_{i,t}}\Bigr)v_{(N),t}>d_{t}^{O}\right].$ (14)
+Taking $\mathbb{E}[\cdot\mid H_{t-1}]$ and using $\text{Pr}(A_{i,t}\mid H_{t-1})=1/N$ yields
+$\mathbb{E}[g_{i,t}\mid H_{t-1}]\ \leq\ \frac{1}{N}\,G(\lambda_{i,t})\ =\ \phi(y_{i,t}).$ (15)
+Fix $\delta>0$ and write $\underline{y}:=y_{\star}-\delta$ and $\underline{\lambda}:=(h^{\prime})^{-1}(\underline{y})$.
+Let $\underline{c}_{\delta}:=-\phi(\underline{y})>0$ (since $\phi(\underline{y})<0$).
+Whenever $y\leq\underline{y}$, monotonicity gives $\phi(y)\leq-\underline{c}_{\delta}$.
+Claim B.2.
+For every bidder $i\in[N]$,
+$\frac{1}{T}\sum_{t=1}^{T}\Pr\big(y_{i,t}\leq y_{\star}-\delta\big)\ \leq\ \frac{(y_{i,1}-y_{\star})^{2}+B^{2}}{2\delta\,\underline{c}_{\delta}}\cdot\frac{1}{\sqrt{T}}.$ (16)
+Proof.
+Let $Z_{t}:=(y_{\star}-y_{i,t})_{+}^{2}$.
+From Equation˜7, $Z_{t+1}\leq(y_{\star}-y_{i,t}+\alpha g_{i,t})^{2}=Z_{t}+2\alpha(y_{\star}-y_{i,t})_{+}g_{i,t}+\alpha^{2}g_{i,t}^{2}$.
+Take $\mathbb{E}[\cdot\mid H_{t-1}]$, use $g_{i,t}^{2}\leq B^{2}$ and Equation˜15:
+$\mathbb{E}[Z_{t+1}\mid H_{t-1}]\leq Z_{t}+2\alpha(y_{\star}-y_{i,t})_{+}\phi(y_{i,t})+\alpha^{2}B^{2}.$
+On event $\{y_{i,t}\leq y_{\star}-\delta\}$ we have $(y_{\star}-y_{i,t})_{+}\geq\delta$ and $\phi(y_{i,t})\leq-\underline{c}_{\delta}$, so the second term on the right hand side is $\leq-2\alpha\delta\underline{c}_{\delta}$.
+Summing over $t$, using $Z_{T+1}\geq 0$ and $\alpha^{2}T=1$, gives Equation˜16.
+∎
+Define “good” event $\mathcal{G}_{t}(\delta):=\bigcap_{i=1}^{N}\{\lambda_{i,t}\geq\underline{\lambda}\}$.
+On $\mathcal{G}_{t}(\delta)$, for every $i$, $\bigl(1+\frac{1}{\lambda_{i,t}}\bigr)\leq\bigl(1+\frac{1}{\underline{\lambda}}\bigr)$, hence $\max_{i}b_{i,t}\leq(1+1/\underline{\lambda})\,v_{(N),t}$.
+Therefore, the total value
+$V_{T}^{I,A}\;\leq\;v_{(N),t}\,\mathbb{I}\left[\Bigl(1+\frac{1}{\underline{\lambda}}\Bigr)v_{(N),t}>d_{t}^{O}\right]\;+\;B\,\mathbb{I}\left[\mathcal{G}_{t}(\delta)^{c}\right].$
+Averaging and taking expectations,
+$\mathbb{E}[V_{T}^{I,A}]\ \leq\ V_{(N)}(\underline{\lambda})\ +\ B\cdot\mathbb{E}\!\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left[\mathcal{G}_{t}(\delta)^{c}\right]\right].$
+By a union bound, $\mathbb{I}\left[\mathcal{G}_{t}(\delta)^{c}\right]\leq\sum_{i=1}^{N}\mathbb{I}\left[\lambda_{i,t}<\underline{\lambda}\right]=\sum_{i=1}^{N}\mathbb{I}\left[y_{i,t}\leq y_{\star}-\delta\right]$, so by Equation˜16, $\mathbb{E}\big[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left[\mathcal{G}_{t}(\delta)^{c}\right]\big]\to 0$ as $T\to\infty$.
+Hence
+$\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\ \leq\ V_{(N)}(\underline{\lambda}).$
+Finally send $\delta\downarrow 0$.
+Since $(h^{\prime})^{-1}$ is continuous, $\underline{\lambda}\uparrow\lambda_{\star}$ and by continuity of $V_{(N)}$, $V_{(N)}(\underline{\lambda})\to V_{(N)}(\lambda_{\star})$.
+Therefore
+$\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\ \leq\ V_{(N)}(\lambda_{\star}).$ (17)
+Step 4: conclude.
+Combining Equations (13) and (17),
+$\liminf_{T\to\infty}\mathbb{E}[V_{T}^{C,A}-V_{T}^{I,A}]\ \geq\ V_{(N)}(\lambda_{\star})-V_{(N)}(\lambda_{\star})=0.$
+This proves the theorem for the case of $\lambda_{\star}\in(0,\infty)$.
+B.2 Boundary case $\lambda_{\star}=0$
+In this section, we consider the boundary case in the proof of Theorem 4.1 where $\lambda_{\star}:=\inf\{\lambda>0:G(\lambda)\geq 0\}=0$.
+Since $G$ is strictly increasing, we have $G(\lambda)>0$ for any $\lambda>0$.
+And because $h^{\prime}$ is strictly increasing, we have for any $y$
+$\phi(y)=\frac{1}{N}G\bigl((h^{\prime})^{-1}(y)\bigr)>0.$ (18)
+Lemma B.2.
+Fix any bidder $i$ in the coordinated bidding algorithm. Let $\alpha:=1/\sqrt{T}$ and define $y_{i,t}:=h^{\prime}(\lambda_{i,t})$. Then for any (negative) $M\in\mathbb{R}$ and any $\eta>0$,
+$\frac{1}{T}\sum_{t=1}^{T}\Pr[y_{i,t}\geq M]\ \leq\ \frac{C_{\eta,B}\,e^{\eta y_{i,1}}}{\eta\,\phi(M)\,e^{\eta M}}\cdot\frac{1}{\sqrt{T}},$ (19)
+where
+$C_{\eta,B}:=1+\frac{\eta^{2}B^{2}}{2}\exp(\eta B)\exp\Bigl(\frac{\eta^{2}B^{2}}{2}\Bigr).$
+In particular, for each fixed $M$,
+$\lim_{T\to\infty}\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}[y_{i,t}\geq M]\Bigr]=0.$
+Proof.
+Fix $M$ and $\eta>0$. Define
+$W_{t}:=\exp(\eta y_{i,t})>0.$
+Using the mirror descent update rule (Equation (7)) $y_{i,t+1}=y_{i,t}-\alpha g_{i,t}$, we have
+$W_{t+1}=W_{t}\exp(-\eta\alpha g_{i,t}).$
+We use the elementary inequality:
+$e^{-x}\leq 1-x+\frac{x^{2}}{2}e^{|x|}.$ (20)
+Apply equation(20) with $x=\eta\alpha g_{i,t}$, noting $|x|\leq\eta\alpha B\leq\eta B$:
+$\exp(-\eta\alpha g_{i,t})-1\leq-\eta\alpha g_{i,t}+\frac{\eta^{2}\alpha^{2}B^{2}}{2}\exp(\eta B).$
+Hence
+$W_{t+1}-W_{t}\leq-\eta\alpha W_{t}g_{i,t}+\frac{\eta^{2}\alpha^{2}B^{2}}{2}\exp(\eta B)\,W_{t}.$
+Taking conditional expectation given $H_{t-1}$ and using $\mathbb{E}[g_{i,t}\mid H_{t-1}]=\phi(y_{i,t})$ (eq.˜8) yields
+$\mathbb{E}[W_{t+1}-W_{t}\mid H_{t-1}]\leq-\eta\alpha W_{t}\phi(y_{i,t})+\frac{\eta^{2}\alpha^{2}B^{2}}{2}\exp(\eta B)\,W_{t}.$ (21)
+Because $\phi$ is increasing and $\phi(y)>0$ for any $y$, on $\{y_{i,t}\geq M\}$ we have $\phi(y_{i,t})\geq\phi(M)$, and on $\{y_{i,t}<M\}$ we have $\phi(y_{i,t})>0$. Therefore
+$W_{t}\phi(y_{i,t})\ \geq\ \phi(M)\,W_{t}\,\mathbb{I}[y_{i,t}\geq M].$
+Taking full expectation in equation (21) gives
+$\mathbb{E}[W_{t+1}-W_{t}]\leq-\eta\alpha\phi(M)\,\mathbb{E}\bigl[W_{t}\mathbb{I}[y_{i,t}\geq M]\bigr]+\frac{\eta^{2}\alpha^{2}B^{2}}{2}\exp(\eta B)\,\mathbb{E}[W_{t}].$ (22)
+Step 1: a uniform bound on $\mathbb{E}[W_{t}]$. Since $|g_{i,t}|\leq B$ and $\mathbb{E}[g_{i,t}\mid\phi_{t-1}]=H(y_{i,t})\geq 0$, Hoeffding’s lemma implies
+$\mathbb{E}\bigl[e^{-\eta\alpha g_{i,t}}\mid H_{t-1}\bigr]\leq\exp\Bigl(\frac{\eta^{2}\alpha^{2}B^{2}}{2}\Bigr).$
+Thus,
+$\mathbb{E}[W_{t+1}\mid H_{t-1}]=W_{t}\mathbb{E}\bigl[e^{-\eta\alpha g_{i,t}}\mid H_{t-1}\bigr]\leq W_{t}\exp\Bigl(\frac{\eta^{2}\alpha^{2}B^{2}}{2}\Bigr),$
+hence iterating and using $\alpha^{2}(t-1)\leq\alpha^{2}T=1$,
+$\sup_{t\leq T+1}\mathbb{E}[W_{t}]\leq\exp(\eta y_{i,1})\exp\Bigl(\frac{\eta^{2}B^{2}}{2}\Bigr)=:K_{\eta,B}\,e^{\eta y_{i,1}}.$ (23)
+Step 2: telescope equation (22). Sum equation (22) over $t=1,\dots,T$ and use $W_{T+1}\geq 0$:
+$\mathbb{E}[W_{1}]\leq\mathbb{E}[W_{T+1}-W_{1}]\leq-\eta\alpha\phi(M)\sum_{t=1}^{T}\mathbb{E}\bigl[W_{t}\mathbb{I}[y_{i,t}\geq M]\bigr]+\frac{\eta^{2}\alpha^{2}B^{2}}{2}\exp(\eta B)\sum_{t=1}^{T}\mathbb{E}[W_{t}].$
+Rearrange and apply equation (23) and $\alpha^{2}T=1$:
+$\eta\alpha\phi(M)\sum_{t=1}^{T}\mathbb{E}\bigl[W_{t}\mathbb{I}[y_{i,t}\geq M]\bigr]\leq\mathbb{E}[W_{1}]+\frac{\eta^{2}B^{2}}{2}\exp(\eta B)\cdot T\cdot K_{\eta,B}e^{\eta y_{i,1}}=e^{\eta y_{i,1}}\cdot C_{\eta,B}.$
+Thus
+$\sum_{t=1}^{T}\mathbb{E}\bigl[W_{t}\mathbb{I}[y_{i,t}\geq M]\bigr]\leq\frac{C_{\eta,B}\,e^{\eta y_{i,1}}}{\eta\alpha\phi(M)}.$ (24)
+Step 3: convert to occupation probability. On $\{y_{i,t}\geq M\}$ we have $W_{t}=e^{\eta y_{i,t}}\geq e^{\eta M}$, hence
+$\mathbb{E}\bigl[W_{t}\mathbb{I}[y_{i,t}\geq M]\bigr]\geq e^{\eta M}\Pr[y_{i,t}\geq M].$
+Summing over $t$ and using equation (24) yields
+$e^{\eta M}\sum_{t=1}^{T}\Pr[y_{i,t}\geq M]\leq\frac{C_{\eta,B}\,e^{\eta y_{i,1}}}{\eta\alpha\phi(M)}.$
+Divide by $T$ and plug $\alpha=1/\sqrt{T}$ to obtain equation (19). ∎
+Lemma B.3.
+Fix any $\delta>0$ and set $M_{\delta}:=h^{\prime}(\delta)\in\mathbb{R}$. Then for every bidder $i$,
+$\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}[\lambda_{i,t}\geq\delta]\Bigr]\leq\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}[y_{i,t}\geq M_{\delta}]\Bigr]\xrightarrow[T\to\infty]{}0.$
+Moreover, by i.i.d. symmetry of values,
+$\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}[\lambda_{i_{t}^{\star},t}\geq\delta]\Bigr]\xrightarrow[T\to\infty]{}0.$
+Proof.
+Since $h^{\prime}$ is increasing, $\lambda_{i,t}\geq\delta\iff y_{i,t}\geq h^{\prime}(\delta)=M_{\delta}$.
+Apply Lemma B.2.
+For the highest-value index $i_{t}^{\star}$, note that conditional on $H_{t-1}$, $\text{Pr}(i_{t}^{\star}=i)=1/N$, hence
+$\mathbb{E}\bigl[\mathbb{I}[\lambda_{i_{t}^{\star},t}\geq\delta]\mid H_{t-1}\bigr]=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[\lambda_{i,t}\geq\delta],$
+and average over $t$ and take expectations. ∎
+Lemma B.4 (Coordinated value converges to $V_{(N)}(0^{+})$).
+Assume the boundary case assumptions above. Then
 […]
-[…]
-Letting ${\varepsilon}=\frac{\Delta T}{2N}$ proves the lemma. ∎
-Lemmas 1 and 2 together prove Theorem 3.1.
-4 Coordination Increases Bidders’ Value
-In this section, we show that coordination can increase the total value of auto-bidders, compared to independent bidding.
-We will focus on a specific auto-bidding algorithm: the mirror-descent algorithm in feng2023online.
-It is the state-of-the-art algorithm for value maximization under RoS constraint, with near-optimal performances for the bidders, achieving $O(\sqrt{T})$ regrets.
-The algorithm maintains a parameter $\lambda_{i,t}$ for each bidder (which corresponds to the Lagrange multiplier of the dual optimization problem), so the bidder overbids by $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$ at each round.
-After observing a utility feedback, the bidder adjusts the parameter $\lambda_{i,t}$ in the opposite direction of the utility (i.e., positive utility decreases $\lambda_{i,t}$).
-The full algorithm is given in Algorithm 3.
-Initialize multiplier $\lambda_{i,1}=1$, learning rate $\alpha=1/\sqrt{T}$;
-for $t=1,2,\dots,T$ do
-   Observe value $v_{i,t}$ and set bid $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$;
-   Observe allocation $x_{i,t}$ and payment $p_{i,t}$, compute utility $g_{i,t}=v_{i,t}\cdot x_{i,t}-p_{i,t}$;
-   Update the multiplier: $\lambda_{i,t+1}=\lambda_{i,t}\exp(-\alpha g_{i,t})$;
-end for
-ALGORITHM 3 Auto-bidding algorithm in feng2023online
-For Algorithm 3, we show that coordinated auto-bidders obtain higher total value than independent auto-bidders. Let
-$V_{t}^{C}=\sum_{i=1}^{N}v_{i,t}x^{C}_{i,t},\quad V_{t}^{I}=\sum_{i=1}^{N}v_{i,t}x^{I}_{i,t}$
-be the total values obtained by the $N$ bidders at round $t$ in the coordinated case and independent case, respectively.
-Theorem 4.1(Coordination Increases Total Value)
-Assume 1, and suppose bidders run Algorithm 3. As $T\to\infty$, in expectation, the bidders’ total value under coordinated bidding is larger than that under independent bidding:
-$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0.$
-To prove Theorem 4.1, we first reduce the $N$-bidder coordinated scenario to a single-bidder problem where the single bidder’s value is the maximum among the coalition members.
-Using this reduction, we prove an interesting fact: under Assumption 1, the multipliers $\lambda_{i,t}^{C}$ of coordinated bidders converge to 0 in the limit as $t\to\infty$, meaning that the coordinated bidders will eventually submit arbitrarily high bids.
-This implies that the coordinated coalition will obtain a higher value than the independent case.
-We present the full proof below.
-4.1 Proof of Theorem 4.1
+Proof.
+In coordination, only $i_{t}^{\star}$ bids against $d_{t}^{O}$, so the realized value in round $t$ is
+$V_{T}^{C,A}=v_{(N),t}\mathbb{I}[\bigl(1+1/\lambda_{i_{t}^{\star},t}\bigr)v_{(N),t}>d_{t}^{O}].$
+Condition on $H_{t-1}$. The fresh randomness $(v_{(N),t},d_{t}^{O})$ is independent of $H_{t-1}$, hence
+$\mathbb{E}[V_{T}^{C,A}\mid H_{t-1}]=V_{(N)}(\lambda_{i_{t}^{\star},t}).$
+Therefore
+$\mathbb{E}[V_{T}^{C,A}]=\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i_{t}^{\star},t})\Bigr].$
+Fix $\varepsilon>0$. By definition of the right-limit $V_{(N)}(0^{+})$ and boundedness $0\leq V_{(N)}(\cdot)\leq B$, there exists $\delta\in(0,1)$ such that
+$|V_{(N)}(\lambda)-V_{(N)}(0^{+})|\leq\varepsilon\qquad\forall\lambda\in(0,\delta].$
+Hence
+$\Bigl|\mathbb{E}[V_{T}^{C,A}]-V_{(N)}(0^{+})\Bigr|\leq\varepsilon+2B\cdot\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}[\lambda_{i_{t}^{\star},t}>\delta]\Bigr].$
+The second term vanishes by Lemma B.3, so $\limsup_{T}|\mathbb{E}[V_{T}^{C,A}]-V_{(N)}(0^{+})|\leq\varepsilon$. Since $\varepsilon$ is arbitrary, the limit is 0. ∎
+Lemma B.5 (Independent value is always $\leq V_{(N)}(0^{+})$).
+For any fixed multipliers $(\lambda_{1},\dots,\lambda_{N})\in(0,\infty)^{N}$, let $V^{I}(\lambda_{1},\dots,\lambda_{N})$ denote the independent-bidding expected total value in one round:
+$V^{I}(\lambda_{1},\dots,\lambda_{N}):=\mathbb{E}\Bigl[\sum_{i=1}^{N}v_{i}\,\mathbb{I}[i\text{ wins under bids }b_{i}=(1+1/\lambda_{i})v_{i}]\Bigr].$
+for all $(\lambda_{1},\dots,\lambda_{N})$,
+$V^{I}(\lambda_{1},\dots,\lambda_{N})\leq V_{(N)}(0^{+}).$
+Consequently, for the dynamic independent process,
+$\mathbb{E}[V_{T}^{I,A}]\leq V_{(N)}(0^{+})\qquad\forall T,\quad\text{and hence}\quad\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\leq V_{(N)}(0^{+}).$
+Proof.
+Fix $(\lambda_{1},\dots,\lambda_{N})$ and pick any $\lambda_{\varepsilon}<\min_{i}\lambda_{i}$.
+Consider the single-bidder scenario where a bidder with value $v_{(N)}:=\max_{i}v_{i}$ bids with multiplier $\lambda_{\varepsilon}$ against the outside bid $d^{O}$ (no internal competition).
+Its bid is $b^{\text{single}}=(1+1/\lambda_{\varepsilon})v_{(N)}$.
+For any realization $(v_{1},\dots,v_{N},d^{O})$, if the coalition loses under independent bidding, then the independent value is 0 and the single-bidder value is $\geq 0$, so the inequality holds.
+If some bidder $i$ wins under independent bidding, then its bid satisfies $b_{i}=(1+1/\lambda_{i})v_{i}>d^{O}$ and $b_{i}\geq\max_{j\neq i}b_{j}$.
+Since $\lambda_{\varepsilon}<\lambda_{i}$, we have $(1+1/\lambda_{\varepsilon})\geq(1+1/\lambda_{i})$, hence
+$b^{\text{single}}=(1+1/\lambda_{\varepsilon})v_{(N)}\geq(1+1/\lambda_{i})v_{i}=b_{i}>d^{O}.$
+Thus the single bidder wins whenever independent wins.
+Moreover, its realized value equals $v_{(N)}\geq v_{i}$, so the realized single-bidder value is pointwise at least the realized independent coalition value.
+Taking expectations yields
+$V_{(N)}(\lambda_{\varepsilon})\ \geq\ V^{I}(\lambda_{1},\dots,\lambda_{N}).$
+Letting $\lambda_{\varepsilon}\downarrow 0$ and using monotone convergence/bounded convergence gives $V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\dots,\lambda_{N})$.
+For the dynamic process, condition on $H_{t-1}$ to fix $(\lambda_{1,t},\dots,\lambda_{N,t})$; then $\mathbb{E}[V_{T}^{I,A}\mid H_{t-1}]=V^{I}(\lambda_{1,t},\dots,\lambda_{N,t})\leq V_{(N)}(0^{+})$.
+Average over $t$ and take expectations.
+∎
+Theorem B.1.
+$\liminf_{T\to\infty}\ \mathbb{E}[V_{T}^{C,A}-V_{T}^{I,A}]\ \geq\ 0.$
+Proof.
+By Lemma B.4, $\lim_{T\to\infty}\mathbb{E}[V_{T}^{C,A}]=V_{(N)}(0^{+})$. By Lemma B.5, $\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\leq V_{(N)}(0^{+})$. Therefore,
+$\liminf_{T\to\infty}\mathbb{E}[V_{T}^{C,A}-V_{T}^{I,A}]\geq\lim_{T\to\infty}\mathbb{E}[V_{T}^{C,A}]-\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\geq V_{(N)}(0^{+})-V_{(N)}(0^{+})=0.$
+B.3 Proof of Theorem 4.2
 Reduction to a single-bidder scenario.
@@
 We show that the $N$-bidder coordinated scenario is “equivalent” to the single-bidder scenario in the following sense:
-Observation 1
+Observation 2.
 $G_{i}^{C}(\lambda_{i})=\frac{1}{N}G(\lambda_{i})$.
-Proof
+Proof.
 In the coordinated scenario, bidder $i$ obtains utility 0 if its value is not the highest among the $N$ bidders.
@@
 Since the $N$-bidder coordinated scenario is equivalent to a single-bidder scenario, we can now analyze the single-bidder scenario.
-We present two useful lemmas regarding the expected utility function $G(\lambda)$ of the single bidder.
-Lemma 3
-$G(\lambda)$ is monotonically increasing in $\lambda$.
-Proof
-See Appendix 0.A. ∎
-Lemma 4
-Under Assumption 1, for any $\lambda>0$, $G(\lambda)\geq\Delta>0$.
-Proof
-By Lemma 3, $G(\lambda)$ is increasing in $\lambda$. So,
+In addition to Lemma B.1, we present the following useful lemma regarding the expected utility function $G(\lambda)$ of the single bidder, under Assumption 1.
+Lemma B.6.
+Under Assumption 1, $G(\lambda)\geq\Delta>0$.
+Proof.
+By Lemma B.1, $G(\lambda)$ is increasing in $\lambda$. So,
 $G(\lambda)$ $\geq\lim_{\mu\to 0^{+}}G(\mu)$
@@
 $=\mathbb{E}\big[(v_{(N)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]$
-$\geq\mathbb{E}\big[(v_{(N-1)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]\geq\Delta>0$
+$\geq\mathbb{E}\big[(v_{(N-1)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]$
+$\geq\Delta>0$
 where the last $\geq$” follows from Assumption 1. ∎
-Intuitively, Lemma 4 shows that the expected utility obtained by the highest-value coalition bidder is always positive (by a margin of $\Delta>0$).
+Intuitively, Lemma B.6 shows that the expected utility obtained by the highest-value coalition bidder is always positive (by a margin of $\Delta>0$).
 This means that the bidder’s multiplier $\lambda_{t}$, following mirror descent updates, will decrease in expectation every round and eventually converge to 0.
@@
 This means that the coalition will eventually submit arbitrarily high bids $b_{i,t}^{C}=(1+\frac{1}{\lambda_{i,t}^{C}})v_{i,t}\to\infty$.
-Lemma 5
-Under Assumption 1, for any $i\in[N],t\in[T]$, with probability at least $1-\exp\big(-\frac{\Delta^{2}t}{32B^{2}N^{2}}\big)$, $\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\big(-\frac{\Delta t}{2N\sqrt{T}}\big)$.
-Proof
-According to Algorithm 3, $\lambda_{i,t+1}^{C}$ is equal to
-$\lambda_{i,t+1}^{C}=\lambda_{i,1}\exp\Big(-\alpha\sum_{t^{\prime}=1}^{t}g_{i,t^{\prime}}\Big)$
+Lemma B.7.
+Under Assumption 1, for any $i\in[N],t\in[T]$, with probability at least $1-\exp\big(-\frac{\Delta^{2}t}{32B^{2}N^{2}}\big)$, $h^{\prime}(\lambda_{i,t+1}^{C})\leq h^{\prime}(\lambda_{i,1})-\frac{\Delta t}{2N\sqrt{T}}$.
+Proof.
+For simplicity, let $y_{i,t}:=h^{\prime}(\lambda_{i,t})$. According to Algorithm 3 and equation (7),
+$y_{i,t+1}=y_{i,t}-\alpha g_{i,t}$ (25)
 where each $g_{i,t^{\prime}}$ is an unbiased estimator of bidder $i$’s expected utility $G_{i}^{C}(\lambda_{i,t^{\prime}})=\mathbb{E}[g_{i,t^{\prime}}]$ at round $t^{\prime}$.
 If bidder $i$ is not the highest-value bidder in round $t^{\prime}$, then $g_{i,t^{\prime}}=0$, meaning that the bidder’s multiplier $\lambda_{i,t^{\prime}}^{C}$ is not updated in that round.
-According to Observation 1 and Lemma 4, bidder $i$’s expected utility is at least
+According to Observation 2 and Lemma B.6, bidder $i$’s expected utility is at least
 $\mathbb{E}[g_{i,t^{\prime}}]=G_{i}^{C}(\lambda_{i,t^{\prime}})=\frac{1}{N}G(\lambda_{i,t^{\prime}})\geq\frac{\Delta}{N}.$
@@
 Let ${\varepsilon}=\frac{t\Delta}{2N}$. We have with probability at least $1-\exp(-\frac{t\Delta^{2}}{32B^{2}N^{2}})$,
-$\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\Big(-\alpha\big(t\cdot\frac{\Delta}{N}-{\varepsilon}\big)\Big)=\lambda_{i,1}\exp\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big).$
+$y_{i,t+1}^{C}\leq y_{i,1}-\alpha\left(t\cdot\frac{\Delta}{N}-{\varepsilon}\right)=y_{i,1}-\frac{t\Delta}{2N\sqrt{T}}.$
 where we plugged in $\alpha=1/\sqrt{T}$. ∎
@@
 Let $V_{(N)}(\lambda)$ be the expected value obtained by a single bidder whose value is the highest value $v_{(N)}$ among $N$ samples from $F$ and who uses multiplier $\lambda>0$ to bid against outside bid $d^{O}\sim D$:
-$V_{(N)}(\lambda)=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Big[v_{(N)}\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
+$V_{(N)}(\lambda)~:=~\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Bigl[v_{(N)}\mathbb{I}\left[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\right]\Bigr]$ (26)
 Because $V_{(N)}(\lambda)$ is monotone in $\lambda$ and bounded, the limit
-$V_{(N)}(0^{+})=\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$
-exists. Note that the expected total value of $N$ coordinated bidders with multipliers $\lambda_{1}^{C},\ldots,\lambda_{N}^{C}$ is
-$V^{C}(\lambda_{1}^{C},\ldots,\lambda_{N}^{C})=\mathbb{E}\Big[v_{i^{*}}\!\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda_{i^{*}}})v_{i^{*}}>d^{O}\big]\Big]=\mathbb{E}\big[V_{(N)}(\lambda_{i^{*}}^{C})\big]$
-where $i^{*}=\operatorname*{arg\,max}_{i\in[N]}v_{i}$. This means that the expected average total value of the coordinated bidders is
-$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{C}\Big]=\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i^{*}_{t},t}^{C})\Big]$ (3)
-where $i^{*}_{t}\in\{1,\dots,N\}$ is the highest-value bidder in round $t$.
-Let $V^{I}(\lambda_{1},\ldots,\lambda_{N})$ be the expected total value of $N$ independent bidders with multipliers $\lambda_{1},\ldots,\lambda_{N}>0$:
-$V^{I}(\lambda_{1},\ldots,\lambda_{N})=\mathbb{E}_{\bm{v}\sim F,d^{O}\sim D}\Big[\sum_{i=1}^{N}v_{i}x_{i}(\bm{b}^{I})\Big]$
-where $\bm{b}^{I}=\big((1+\tfrac{1}{\lambda_{1}})v_{1},\ldots,(1+\tfrac{1}{\lambda_{N}})v_{N},d^{O}\big)$.
-Lemma 6
-$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$, for any $\lambda_{1},\ldots,\lambda_{N}>0$.
-Proof
-Pick $\lambda_{\varepsilon}<\min\{\lambda_{1},\dots,\lambda_{N}\}$.
-If some bidder $i\in[N]$ wins the item under independent bidding, then their bid $b_{i}^{I}=(1+\frac{1}{\lambda_{i}})v_{i}$ is larger than the outside bid $d^{O}$.
-In that case, suppose the highest-value bidder uses multiplier $\lambda_{\varepsilon}$, then their bid is $(1+\frac{1}{\lambda_{\varepsilon}})v_{(N)}\geq(1+\frac{1}{\lambda_{i}})v_{i}=b_{i}^{I}>d^{O}$, so the highest-value bidder also wins in the single-bidder scenario.
-Note that the value obtained by the highest-value bidder is at least $v_{i}$, so the expected value $V_{(N)}(\lambda_{\varepsilon})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$.
-Letting $\lambda_{\varepsilon}\to 0$ proves the lemma.
+$V_{(N)}(0^{+})~:=~\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$
+exists. Note that the expected total value of the $N$ coordinated bidders with multipliers $\lambda_{1}^{C},\ldots,\lambda_{N}^{C}$ is
+$V^{C}(\lambda_{1}^{C},\ldots,\lambda_{N}^{C})~=~\mathbb{E}\Big[v_{i^{*}}\mathbb{I}\big[(1+\tfrac{1}{\lambda_{i^{*}}})v_{i^{*}}>d^{O}\big]\Big]~=~\mathbb{E}\big[V_{(N)}(\lambda_{i^{*}}^{C})\big]$
+where $i^{*}=\operatorname*{arg\,max}_{i\in[N]}v_{i}$. This means that the expected average total value of the coordinated bidders during $T$ rounds is
+$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{C,A}\Big]~=~\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i^{*}_{t},t}^{C})\Big]$ (27)
+where $i^{*}_{t}\in[N]$ is the highest-value bidder in round $t$.
+Lemma B.8.
+For any coordination mechanism $\mathcal{G}$ and any round $t$,
+$\mathbb{E}\!\left[V_{t}^{\mathcal{G}}\right]~\leq~V_{(N)}(0^{+})~=~\mathbb{E}_{v_{(N)}\sim F_{(N)}}\big[v_{(N)}\big].$
+Consequently, for every $T\geq 1$,
+$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{\mathcal{G}}\Big]~\leq~V_{(N)}(0^{+}).$
+Proof.
+Fix a round $t$ and write $v_{(N),t}\coloneqq\max_{i\in[N]}v_{i,t}$.
+In a single-item auction, the coalition’s realized value is either 0 (if the outside bid wins) or equals the realized value of the winning coalition bidder, which is always at most $v_{(N),t}$.
+Hence $V_{t}^{\mathcal{G}}\leq v_{(N),t}$ almost surely, and taking expectations gives $\mathbb{E}[V_{t}^{\mathcal{G}}]\leq\mathbb{E}[v_{(N),t}]$.
+Finally, since bids and $d^{O}$ are bounded, $(1+\tfrac{1}{\lambda})v_{(N)}\to+\infty$ as $\lambda\downarrow 0$, so the single bidder wins with probability $1$ in the limit and thus $V_{(N)}(0^{+})=\mathbb{E}[v_{(N)}]$.
 ∎
-Proof(Proof of Theorem 4.1)
-By Lemma 5, with constant $c\coloneqq\frac{\Delta^{2}}{32B^{2}N^{2}}$, for any fixed $i$ and $t\geq 1$, the coordinated multiplier $\lambda_{i,t+1}^{C}$ satisfies
-$\Pr\!\left[\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\right]\;\geq\;1-e^{-ct}.$ (4)
+Proof of Theorem 4.2.
+By Lemma B.7, with constant $c\coloneqq\frac{\Delta^{2}}{32B^{2}N^{2}}$, for any fixed $i$ and $t\geq 1$, the coordinated multiplier $\lambda_{i,t+1}^{C}$ satisfies
+$\Pr\!\left[h^{\prime}(\lambda_{i,t+1}^{C})\leq h^{\prime}(\lambda_{i,1})-\frac{t\Delta}{2N\sqrt{T}}\right]\;\geq\;1-e^{-ct}.$ (28)
 Fix $\varepsilon>0$. Since $V_{(N)}(0^{+})$ exists, there is $\delta=\delta(\varepsilon)>0$ such that
 $\big|V_{(N)}(0^{+})-V_{(N)}(\lambda)\big|<\varepsilon\quad\text{for all }\lambda\in(0,\delta].$ (29)
+Since $h^{\prime}$ is strictly increasing and $\lim_{\lambda\downarrow 0}h^{\prime}(\lambda)=$-$\infty$, we can define $b(\delta):=(h^{\prime})^{-1}(\delta)$.
 Define
-$m_{T}\;\coloneqq\;\left\lceil\frac{2N\sqrt{T}}{\Delta}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\right\rceil~=~O\Big(\sqrt{T}\log\frac{1}{\delta}\Big),$
+$m_{T}\;\coloneqq\;\left\lceil\frac{2N\sqrt{T}}{\Delta}(h^{\prime}(\lambda_{i,1})-b(\delta))\right\rceil~=~O\Big(\sqrt{T}(-b(\delta))\Big),$
 so for all $t\geq m_{T}$,
-$\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\leq\lambda_{i,1}\exp\!\Big(-\frac{m_{T}\Delta}{2N\sqrt{T}}\Big)\leq\delta.$
+$h^{\prime}(\lambda_{i,1})-\frac{t\Delta}{2N\sqrt{T}}\leq h^{\prime}(\lambda_{i,1})-\frac{m_{T}\Delta}{2N\sqrt{T}}\leq b(\delta).$
 Let $E_{i,t}$ be the event inside the probability in (28). A union bound over $i=1,\dots,N$ and $t\geq m_{T}$ gives
 $\Pr\!\Big[\bigcap_{i=1}^{N}\bigcap_{t=m_{T}}^{T}E_{i,t}\Big]\;\geq\;1-NTe^{-cm_{T}}\;\eqqcolon\;1-p_{T}.$ (30)
-On this event, for all $t\geq m_{T}$ we have $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by Eq. (5),
+On this event, for all $t\geq m_{T}$ we have $h^{\prime}(\lambda_{i_{t}^{*},t}^{C})\leq b(\delta)$, which is equivalent to $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by Eq. (29),
 $\big|V_{(N)}(0^{+})-V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\big|<\varepsilon.$
@@
 Taking expectations and using that the per-round gap is $\leq B$ on the complement of the event in (30), we have
-$\mathbb{E}\bigg[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\bigg]\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon+B\,p_{T}$
-$\;\leq\;\frac{O(\sqrt{T}\log\tfrac{1}{\delta})}{T}\,B+\varepsilon+BNTe^{-cO(\sqrt{T}\log\tfrac{1}{\delta})}~\leq~3{\varepsilon}$
+$\mathbb{E}\bigg[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\bigg]$
+$\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon+B\,p_{T}$
+$\;\leq\;\frac{O(\sqrt{T}(-b(\delta)))}{T}\,B+\varepsilon+BNTe^{-cO(\sqrt{T}(-b(\delta)))}~\leq~3{\varepsilon}$
 for sufficiently large $T$. So we obtain
 $\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]\,=\,V_{(N)}(0^{+}).$
-Combining with Lemma 6, we conclude that
-$V_{(N)}(0^{+})\;\geq\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{I}(\lambda_{1,t}^{I},\ldots,\lambda_{N,t}^{I})\Big]\;=\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{I}\Big],$
+Combining with Lemma B.8, we conclude that
+$V_{(N)}(0^{+})\;\geq\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{\mathcal{G}}\Big],$
 which proves Theorem 4.2. ∎
-5 Extension to Non-i.i.d. Valuations
-We relax the i.i.d.
-assumption and allow bidder-specific value distributions.
-In each round $t\in[T]$, bidder $i\in[N]$ draws a value $v_{i,t}\sim F_{i}$ (independently across bidders and rounds), and the outside bid $d^{O}_{t}$ is drawn from $D$.
-Values and outside bids are bounded by $B>0$.
-Let $V_{t}^{C}$ and $V_{t}^{I}$ denote the coalition’s total value achieved in round $t$ under coordinated and independent bidding, respectively.
-Let $U_{i}^{\cdot,A}$ denote the cumulative RoS violation (achieved value minus payment) of bidder $i$ over $T$ rounds under scenario ${\cdot}\in\{C,I\}$ when all bidders follow auto-bidding algorithm $A$.
-Total RoS violation.
-Our first result shows that, in the non-i.i.d. setting, coordination strictly reduces the total RoS constraint violation, namely, the total utility of the coalition bidders is improved:
-Theorem 5.1(Coordination reduces Total RoS Violation, non-iid case)
-Assume 1.
-For any auto-bidding algorithm $A$ that (weakly) overbids, with probability at least $1-\exp\!\big(-\tfrac{\Delta^{2}T}{32B^{2}}\big)$, the total utility $\sum_{i=1}^{N}U_{i}^{C,A}\geq\sum_{i=1}^{N}U_{i}^{I,A}+\frac{\Delta T}{2}$.
-See Appendix 0.B for the proof.
-Total value.
-To compare total value, we require a per-bidder margin condition that ensures each bidder wins with a positive average advantage when they are the highest-value coalition member.
-For bidder $i$, define $\Delta_{i}:=\mathbb{E}\left[(v_{i}-d^{O})\,\mathbb{I}\{v_{i}=\max_{j\in[N]}v_{j}\}\right]$.
-Assumption 2 (Positive per-bidder winning margin)
-For every bidder $i\in[N]$, $\Delta_{i}>0$.
-Theorem 5.2(Coordination Increases Total Value, non-iid case)
-Assume 2, and suppose bidders run Algorithm 3. Then, we have $\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0$.
-See Appendix 0.C for the proof.
-The findings in this section suggest that coordination improves upon independent bidding even in the presence of asymmetry inside the coalition, demonstrating the robustness of our conclusions.
-Figure Setting NN $T$ Total Utility (I) Total Utility (C) Total Value (I) Total Value (C)
-1(a) i.i.d. 2 4000 -0.011 0.220 0.643 0.666
-1(b) i.i.d. 4 4000 -0.077 0.302 0.773 0.800
-1(c) i.i.d. 3 4000 -0.048 0.154 0.713 0.749
-2(a) non-i.i.d. 2 10000 0.049 0.219 0.715 0.718
-2(b) non-i.i.d. 3 20000 -0.014 0.258 0.620 0.633
-2(c) non-i.i.d. 5 20000 -0.062 0.619 0.814 0.819
-3(a) Real data 4 20000 -0.040 0.156 $\pm$ 0.012 0.614 $\pm$ 0.016 0.929
-3(b) Real data 5 20000 -0.065 0.174 $\pm$ 0.012 0.607 $\pm$ 0.012 0.959
-Table 1: Performance summary. Confidence intervals below $\pm 0.002$ are omitted.
-(a) $N=2,F=U[0,1],D=U[0,0.9]$
-(b) $N=4,F=U[0,1],D=U[0,1]$
-(c) $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$
-Figure 1: Experiments under i.i.d auto-bidders.
-6 Experiments
-To complement our theoretical analysis, we conduct a series of experiments evaluating the performance of coordinated auto-bidding, on both synthetic data and a public real-world dataset (iPinYou [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou]).
-Unless otherwise stated, all metrics are averaged over 100 independent simulations; shaded regions in figures depict 95% confidence intervals computed as $\bar{x}\pm 1.96\cdot s$, where $\bar{x}$ and $s$ are the sample mean and standard deviation over the 100 simulations.
-6.1 Experiment Setup
-For all experiments, we run Algorithm 3 with learning rate $\alpha=1/\sqrt{T}$, where $T$ is the total number of auction rounds.
-The horizon $T$ may vary across experiments because different settings exhibit different convergence rates.
-The values of $T$ and $N$ (number of coalition bidders) used in each setting are summarized in Table 1.
-6.1.1 Synthetic Data
-Symmetric setting.
-Each coalition bidder $i$ draws values i.i.d.
-from uniform distribution $U[0,1]$.
-The outside bid is sampled from $D=U[0,0.9]$ (Figure 1(a)), $D=U[0,1]$ (Figure 1(b)), or $D=\mathrm{Beta}(3,2)$ (Figure 1(c)), representing different market conditions faced by the coalition bidders.
-Asymmetric setting.
-To stress-test the mechanism under heterogeneity, we sample each bidder’s values from distinct families (Uniform, Beta, or truncated Gaussians with bidder-specific parameters), as summarized in Figure 2.
-The outside-bid distribution $D$ varies with the scenario: $U[0.2,0.8]$ in Figure 2(a), $\mathrm{Beta}(3,5)$ in Figure 2(b), and $\mathrm{Beta}(2,8)$ in Figure 2(c).
-Asymmetries arise from both inside the coalition and the outside market conditions.
-Full distributional specifications for the synthetic experiments are provided in the figure captions (Figures 1 and 2). All distributions conform to Assumption 1.
-6.1.2 Real-World Data
-We also experimented with the public iPinYou dataset [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou], released for a real-time bidding competition organized by iPinYou.
-We use the Season 2 test split, which contains 2,521,6302{,}521{,}630 auction records from $5$ advertisers.
-Each record provides the bidding (winning) price and the corresponding advertiser ID.
-The real-world data impose three limitations: (i) only the winning price is observed, (ii) bidder identities are unavailable, and (iii) there is no label indicating whether any bidders coordinate.
-To obtain bidder valuations consistent with the data while abstracting away unobserved identities and strategies, we adopt an empirical i.i.d.
-model: we aggregate all observed winning prices into a common pool and, in each round, draw each coalition bidder’s value independently and uniformly at random with replacement from this pool.
-This is equivalent to sampling i.i.d.
-values from the empirical price distribution induced by the dataset.
-Following prior work [chen_coordinated_2023], we normalize all prices to $[0,1]$ for numerical stability.
-The outside bidder’s bid is drawn from the same empirical distribution and then scaled by a random factor from $U[1,2]$ to maintain competitive pressure.
-We report results for coalitions with $N\in\{4,5\}$ bidders under this setup.
-6.2 Results
-Metrics.
-Across all experiments, we evaluate three metrics for both Independent (Algorithm 1) and Coordination (Algorithm 2): (a) utility (achieved value minus payment), (b) achieved value, and (c) multiplier.
-Higher utility indicates lower RoS violation; in particular, utility $\geq 0$ implies that the RoS constraint is satisfied.
-Figures report per-bidder trajectories, whereas Table 1 aggregates to coalition-level totals (sum across bidders).
-All statistics are averaged over 100 independent runs.
-Plot notation.
-In the figures, labels of the form bidder $i$ -- I denote bidder $i$ under the Independent scenario, whereas bidder $i$ -- C denotes the Coordination scenario.
-Shaded regions show 95% confidence intervals over 100 simulations.
-Results on synthetic data
-In the symmetric settings (Figure 1), coordination sharply increases both each bidder’s utility (hence reducing RoS violation) and the coalition’s total achieved value relative to independent bidding.
-Although our theory guarantees value improvement at the coalition (total) level, the experiments further show that every individual bidder benefits from coordination by obtaining higher value.
-Moreover, all multipliers converge to 0, consistent with our theory, and per-bidder statistics (violation, value, and multiplier trajectories) closely track one another due to the i.i.d.
-value draws.
-In the asymmetric settings (Figure 2), the aggregate gains largely persist: coordination improves total value on average and typically improves violations.
-However, heterogeneity induces dispersion in learning dynamics – bidders with different value distributions exhibit different convergence rates of the multiplier, and individual value improvements are not guaranteed for every bidder even though the coalition-level total value increases.
-In cases where each bidder has a non-negligible chance to be the winning bidder, we also observe that the multipliers tend to converge toward 0 empirically.
-Results on real-world data.
-Consistent with the synthetic data experiments, real-world data results show improvements in both individual RoS violation and achieved value, accompanied by multipliers converging steadily toward 0.
-7 Discussion
-In this paper, we studied coordinated auto-bidding mechanisms for value maximization under RoS constraints, to understand how cooperation among learning agents enhances efficiency in repeated auctions.
-We proposed a simple yet effective coordination protocol and theoretically proved that coordination improves the coalition’s total value, as well as RoS constraint compliance, under mild assumptions.
-Our empirical results further validate these findings, showing higher achieved value and consistent convergence patterns.
-The demonstrated benefit of coordination over independent bidding might provide useful guidance for industrial practice.
-Building on these findings, we outline several directions for future work:
-  • Auto-bidding outside bidders: While we assumed i.i.d.
-outside bids (a reasonable assumption when outside bidders are uniformly sampled from a population every round), a natural extension is to incorporate auto-bidding outside bidders, allowing both coalition members and external participants to adjust their bids dynamically.
-  • Alternative coordination mechanisms: Another direction is to explore alternative coordination mechanisms, such as highest-bid coordination, to examine how they would affect the bidders’ overall welfare.
-Highest-bid coordination is analogous to highest-value coordination in the symmetric setting, but not in the asymmetric setting.
-  • Other auction formats: Finally, it would be valuable to extend our analysis to other auction formats, especially non-truthful auctions such as first-price and generalized second-price auctions.
-Such extensions would help test the robustness of our theoretical insights and clarify how coordination interacts with strategic bidding incentives in broader market environments.
-Appendix 0.A Proof of Lemma 3
-Write the single-bidder’s expected utility as $G(\lambda)=\mathbb{E}_{v\sim F_{(N)}}G(\lambda;v)$ where $G(\lambda;v)$ is the bidder’s expected utility conditioning on having value $v$, which is equal to
-$G(\lambda;v)$ $=\mathbb{E}_{d^{O}\sim D}\big[(v-d^{O})\cdot\mathbb{I}[(1+\tfrac{1}{\lambda})v>d^{O}]\big]$
-$=vx\big((1+\tfrac{1}{\lambda})v\big)-p\big((1+\tfrac{1}{\lambda})v\big),$
-where $x(b)$ is the winning probability when the bidder bids $b$ and $p(b)$ is the expected payment when the bidder bids $b$.
-Since the second-price auction is a truthful auction, by Myerson’s payment identity for truthful auctions, we have
-$p(b)=xb-\int_{0}^{b}x(z)\mathrm{d}z.$ (7)
-So,
-$G(\lambda;v)$
-$=vx\big((1+\tfrac{1}{\lambda})v\big)-(1+\tfrac{1}{\lambda})vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz$
-$=-\tfrac{1}{\lambda}vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz.$
-Take derivative with respect to $\lambda$:
-$\tfrac{\partial G(\lambda;v)}{\partial\lambda}$
-$=\tfrac{1}{\lambda^{2}}vx\big((1+\tfrac{1}{\lambda})v\big)-\tfrac{1}{\lambda}vx^{\prime}\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})+x\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})$
-$=\tfrac{v^{2}}{\lambda^{3}}x^{\prime}\big((1+\tfrac{1}{\lambda})v\big)$
-$\geq 0$
-where the last inequality is because $x^{\prime}(b)\geq 0$. Integrating over $v\sim F_{(N)}$, we obtain $G^{\prime}(\lambda)\geq 0$, hence $G(\lambda)$ is increasing.
-Appendix 0.B Proof of Theorem 5.1
+Appendix C Omitted Proofs in Section 5
+C.1 Proof of Theorem 5.1
 The argument parallels the proof of Theorem 3.1.
 Let $i^{*}_{t}=\arg\max_{i\in[N]}v_{i,t}$ denote the index of the highest-value bidder in round $t$.
-Conditioned on the history $H_{t-1}$, we analyze the expected increment $\Delta_{i,t}=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$ for bidder $i$:
-$\mathbb{E}_{\bm{v}_{t}\sim F,\,d_{t}^{O}\sim D}\!\left[\Delta_{i,t}\mid H_{t-1}\right]$
-$=\mathbb{E}_{\bm{v}_{t},\,d_{t}^{O}}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\Delta_{i,t}\mid H_{t-1}\right]$
-$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
-$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
-$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
-$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
-$\geq\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\,\big(\max_{j\neq i}v_{j,t}-d_{t}^{O}\big)_{+}\right]$
-$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}]\cdot(v_{i,t}-d_{t}^{O})\right].$
+Conditioned on the history $H_{t-1}$, we analyze the expected difference $\Delta_{i,t}=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$ for bidder $i$:
+$\mathbb{E}_{\bm{v}_{t}\sim F,\,d_{t}^{O}\sim D}\big[\Delta_{i,t}\mid H_{t-1}\big]$
+$=\mathbb{E}_{\bm{v}_{t},\,d_{t}^{O}}\big[\mathbb{I}[i^{*}_{t}=i]\cdot\Delta_{i,t}\mid H_{t-1}\big]$
+$=\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\Big]+\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\Big]$
+$=\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\Big]+\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\Big]$
+$\geq\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\,\big(\max_{j\neq i}v_{j,t}-d_{t}^{O}\big)_{+}\Big]+\mathbb{E}\Big[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}]\cdot(v_{i,t}-d_{t}^{O})\Big].$
 Summing over all bidders $i\in[N]$, the indicator $\mathbb{I}[i^{*}_{t}=i]$ ensures that only the highest-value bidder contributes in each round. Hence,
@@
 $\geq\Delta,$
-where the second inequality follows from Assumption 1. Finally, applying Azuma’s inequality (in a similar way as in the proof of Theorem 3.1) completes the proof.
-Appendix 0.C Proof of Theorem 5.2
+where the second inequality follows from Assumption 1. Summing over $t\in[T]$ gives $\mathbb{E}[\sum_{i=1}^{N}U_{i,T}^{C,A}-\sum_{i=1}^{N}U_{i,T}^{I,A}]\geq\Delta T$.
+C.2 Proof of Theorem 5.2
 The argument parallels the proof of Theorem 4.2.
-Let $G_{i}^{C}(\lambda_{i})$ denote bidder $i$’s expected utility when $v_{i}\sim F_{i}$ and the coalition uses multiplier $\lambda_{i}>0$:
+Let $G_{i}^{C}(\lambda_{i})$ denote bidder $i$’s expected utility when $v_{i}\sim F_{i}$ and bidder $i$ uses multiplier $\lambda_{i}>0$ in the coordinated scenario:
 $G_{i}^{C}(\lambda_{i})\;=\;\mathbb{E}_{v_{1}\sim F_{1},\dots,v_{N}\sim F_{N},\;d^{O}\sim D}\big[(v_{i}-d^{O})\,x_{i}(\bm{v},\lambda_{i})\big],$
 where $x_{i}(\bm{v},\lambda_{i})=1$ if $v_{i}=\max_{j\in[N]}\{v_{j}\}$ and $(1+\tfrac{1}{\lambda_{i}})v_{i}>d^{O}$, and 0 otherwise.
-By reusing the proof of Lemma 3 (Appendix 0.A) with $D$ replaced by the distribution of $\max\{d^{O},\max_{j\neq i}v_{j}\}$, we claim that $G_{i}^{C}(\lambda_{i})$ is increasing in $\lambda_{i}$.
-Hence,
-$G_{i}^{C}(\lambda_{i})\geq\lim_{\mu\to 0^{+}}G_{i}^{C}(\mu)=\mathbb{E}\!\left[(v_{i}-d^{O})\,\mathbb{I}\!\left\{v_{i}=\max_{j\in[N]}\{v_{j}\}\right\}\right]=:\Delta_{i},$
+By reusing the proof of Lemma B.1 with $D$ replaced by the distribution of $\max\{d^{O},\max_{j\neq i}v_{j}\}$, we claim that $G_{i}^{C}(\lambda_{i})$ is increasing in $\lambda_{i}$. Hence,
+$G_{i}^{C}(\lambda_{i})~\geq~\lim_{\mu\to 0^{+}}G_{i}^{C}(\mu)~=~E\Big[(v_{i}-d^{O})\,\mathbb{I}\big\{v_{i}=\max_{j\in[N]}\{v_{j}\}\big\}\Big]~=:~\Delta_{i},$
 where the limit exists because $G_{i}^{C}(\cdot)$ is bounded and monotone.
-We now present the parallel version of Lemma 5.
-Lemma 7
+We now present the parallel version of Lemma B.7.
+Lemma C.1.
 Under Assumption 2, for any $i\in[N]$ and $t\geq 1$, with probability at least $1-\exp\!\big(-\tfrac{\Delta_{i}^{2}t}{32B^{2}}\big)$,
-$\lambda_{i,t+1}^{C}\;\leq\;\lambda_{i,1}\,\exp\!\Big(-\frac{\Delta_{i}\,t}{2\sqrt{T}}\Big).$
-Proof
+$h^{\prime}(\lambda_{i,t+1}^{C})\leq h^{\prime}(\lambda_{i,1})-\frac{\Delta_{i}t}{2\sqrt{T}}.$
+Proof.
 The expected utility $\mathbb{E}[g_{i,t^{\prime}}]$ is lower bounded by
 $\mathbb{E}[g_{i,t^{\prime}}]\;=\;G_{i}^{C}(\lambda_{i,t^{\prime}})\;\geq\;\Delta_{i}.$
-Applying Azuma’s inequality (exactly as in the proof of Lemma 5) completes the argument. ∎
+Applying Azuma’s inequality (exactly as in the proof of Lemma B.7) completes the argument. ∎
 Write $v_{(N)}\coloneqq\max_{j\in[N]}v_{j}$ with $v_{j}\sim F_{j}$ mutually independent. Its CDF is $F^{\max}(x)=\Pr[v_{(N)}\leq x]=\prod_{j=1}^{N}F_{j}(x)$. For $\lambda>0$, define
-$V_{(N)}(\lambda)$ $\;\coloneqq\;\mathbb{E}_{\,v_{(N)}\sim F^{\max},\,d^{O}\sim D}\Big[\,v_{(N)}\cdot\mathbb{I}\big\{(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big\}\Big],$
-$V_{(N)}(0^{+})$ $\;\coloneqq\;\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda).$
+$V_{(N)}(\lambda)$ $~\coloneqq~\mathbb{E}_{\,v_{(N)}\sim F^{\max},\,d^{O}\sim D}\Big[\,v_{(N)}\cdot\mathbb{I}\big\{(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big\}\Big],$
+$V_{(N)}(0^{+})$ $~\coloneqq~\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)~=~\mathbb{E}_{\,v_{(N)}\sim F^{\max}}\big[\,v_{(N)}\big].$
 The limit $V_{(N)}(0^{+})=\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$ exists because $V_{(N)}(\lambda)$ is monotone and bounded.
@@
 special case $F_{1}=\cdots=F_{N}=F$, $F^{\max}$ coincides with the order-statistic distribution $F_{(N)}$, recovering the original definition.
-In the non-i.i.d. case, Lemma 6 still holds:
-Lemma 8(copy of Lemma 6)
-$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$.
-Proof(Proof of Theorem 5.2)
+In the non-i.i.d. case, Lemma B.8 still holds:
+Lemma C.2 (copy of Lemma B.8).
+For any coordination mechanism $\mathcal{G}$ and any round $t$,
+$\mathbb{E}\!\left[V_{t}^{\mathcal{G}}\right]~\leq~V_{(N)}(0^{+})~=~\mathbb{E}_{v_{(N)}\sim F^{\max}}\big[v_{(N)}\big].$
+Consequently, for every $T\geq 1$,
+$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{\mathcal{G}}\Big]~\leq~V_{(N)}(0^{+}).$
+Proof of Theorem 5.2.
 Fix $\varepsilon>0$ and choose $\delta=\delta(\varepsilon)>0$ as in Equation (29). Let
-$\Delta_{\min}\coloneqq\min_{i\in[N]}\Delta_{i}>0,\qquad c_{\min}\coloneqq\frac{\Delta_{\min}^{2}}{32B^{2}}$
-Define
-$m_{T}\;\coloneqq\;\Big\lceil\frac{2\sqrt{T}}{\Delta_{\min}}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\Big\rceil.$
-By Lemma 7, for each $i$ and every $t\geq m_{T}$, $\lambda_{i,t+1}^{C}\leq\delta$ holds with probability at least $1-\exp(-c_{\min}t)$. A union bound over $i=1,\dots,N$ and $t=m_{T},\dots,T$ yields
+$\Delta_{\min}\coloneqq\min_{i\in[N]}\Delta_{i}>0\qquad c_{\min}\coloneqq\frac{\Delta_{\min}^{2}}{32B^{2}}>0.$
+Define $b(\delta):=(h^{\prime})^{-1}(\delta)$ and
+$m_{T}\;\coloneqq\;\Big\lceil\frac{2\sqrt{T}}{\Delta_{\min}}(h^{\prime}(\lambda_{i,1})-b(\delta))\Big\rceil.$
+By Lemma C.1, for each $i$ and every $t\geq m_{T}$, $\lambda_{i,t+1}^{C}\leq\delta$ holds with probability at least $1-\exp(-c_{\min}t)$. A union bound over $i=1,\dots,N$ and $t=m_{T},\dots,T$ yields
 $\Pr\!\Big[\bigcap_{i=1}^{N}\bigcap_{t=m_{T}}^{T}\{\lambda_{i,t+1}^{C}\leq\delta\}\Big]\;\geq\;1-NT\,e^{-c_{\min}m_{T}}\;\eqqcolon\;1-p_{T}.$
@@
 on the above high-probability event; on its complement the gap is $\leq B$. Taking expectations and using $p_{T}\leq NTe^{-c_{\min}m_{T}}$ gives
-$\mathbb{E}\Big[\Big|V_{(N)}(0^{+})-\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\Big]\;\leq\;\frac{m_{T}}{T}B+\varepsilon+B\,p_{T}\;\xrightarrow[T\to\infty]{}\;0.$
+$\mathbb{E}\Big[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\Big]\;\leq\;\frac{m_{T}}{T}B+\varepsilon+B\,p_{T}\;\xrightarrow[T\to\infty]{}\;0.$
 Hence
-$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]=V_{(N)}(0^{+}).$
-Finally, combining with Lemma 8 yields the desired inequality against the independent-bidding benchmark, completing the proof. ∎
-Appendix 0.D Additional Experiment Results
-This appendix presents additional figures in the experiments: Figure 2 for the asymmetric setting on synthetic data; Figure 3 for the real-world dataset.
-(a) $F_{1}=\mathrm{Beta}(2,5),F_{2}=\mathrm{Beta}(5,2),D=U[0.2,0.8],N=2$
-(b) $F_{1}=\mathrm{Beta}(2,6),F_{2}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.6,0.15),F_{3}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.4,0.2),D=\mathrm{Beta}(3,5),N=3$
-(c) $F_{1}=U[0.2,0.8],F_{2}=\mathrm{Beta}(4,3),F_{3}=\mathrm{Beta}(6,2),F_{4}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.5,0.1),F_{5}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.7,0.12),D=\mathrm{Beta}(2,8),N=5$
+$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]~=~V_{(N)}(0^{+}).$
+Finally, combining with Lemma C.2 yields the desired inequality against the independent-bidding total value, completing the proof. ∎
+Appendix D Additional Experiment Results
+This appendix presents additional figures in the experiments: Figure 3 for the symmetric setting on synthetic data; Figure 4 for the asymmetric setting on synthetic data; Figure 5 for the real-world dataset.
+(a) $N=4,F=U[0,1],D=U[0,1]$
+(b) $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$
+Figure 3: Experiments under i.i.d auto-bidders.
+(a) $N=3,F_{1}=\mathrm{Beta}(2,6),F_{2}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.6,0.15),F_{3}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.4,0.2),D=\mathrm{Beta}(3,5)$
+(b) $N=5,F_{1}=U[0.2,0.8],F_{2}=\mathrm{Beta}(4,3),F_{3}=\mathrm{Beta}(6,2),F_{4}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.5,0.1),F_{5}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.7,0.12),D=\mathrm{Beta}(2,8)$
 Figure 4: Experiments under non-i.i.d auto-bidders.
-(a) $N=4$
-(b) $N=5$
-Figure 3: Experiments in real-world datasets.
+Figure 5: Experiments in real-world datasets. ($N=5$).